### 09 Mar 2026 v1

# Theorem of the heart for Weibel’s homotopy -theory

Abstract.
In this paper we prove the theorem of the heart for Weibel’s homotopy -theory Namely, if is a small stable -category with a bounded -structure, then the realization functor induces an equivalence of spectra In a certain sense this result is dual to the Dundas-Goodwillie-McCarthy theorem. We deduce the dévissage theorem for of abelian categories, also on the level of spectra (in all degrees). More generally, we prove these results for dualizable categories with nice -structures and for the so-called coherently assembled abelian categories.
The proof is heavily based on another new result, which is a much stronger version of Barwick’s theorem of the heart. Its special case states the following: if is a small stable category with a bounded -structure, such that for some the realization functor induces isomorphisms on between the objects of then the map is an isomorphism for and a monomorphism for Moreover, we prove that these estimates are sharp, even for dg categories over a field. In particular, the naive -theoretic theorem of the heart fails for
Contents
- 0 Introduction
1 Coherently assembled abelian categories and exact categories
- 1.1 Reminder on exact categories
- 1.2 Accessibly exact categories
- 1.3 Grothendieck exact categories
- 1.4 Criterion of -accessibility for exact categories
- 1.5 Presentable stable envelopes
- 1.6 Reminder on compactly assembled and dualizable categories
- 1.7 Coherently assembled exact categories
- 1.8 Coherently assembled abelian categories
- 1.9 The situation of dévissage
- 2 Compactly assembled -structures on dualizable categories
- 3 Refined -theoretic theorem of the heart for dualizable categories
- 4 Coconnectivity estimates
- 5 Theorem of the heart for
- 6 Dévissage theorems for -theory and homotopy -theory
- 7 Sharpness of the estimates
- 8 Examples of compactly assembled -structures and coherently assembled abelian categories
- References
0. Introduction
In this paper we study the non-connective -theory and Weibel’s homotopy -theory of small stable -categories with bounded -structures, and more generally for dualizable categories with suitably defined nice -structures. In the latter case we consider the invariants and as defined in [E24]. To fix the ideas, in the introduction we give the formulations of our main results in the context of small categories, mentioning the more general versions for dualizable categories and the so-called coherently assembled abelian categories (Definition 1.41). For brevity we will say “stable category” instead of “stable -category”.
Originally, Weibel’s definition of was given in the -linear context, see [Wei89, Wei13, Tab15]. The natural generalization for stable categories is explained for example in [E25c, Section 8, Corollary 8.3]. For a small stable category the spectrum is given by the geometric realization
where for we put
For a small abelian category we put
Our main result is the following theorem, which is a special case of Theorem 5.1.
Theorem 0.1.
Let be a small stable category with a bounded -structure Then the realization functor induces an equivalence of spectra
This theorem can be considered as a dual version of the Dundas-Goodwillie-McCarthy theorem [DGM13, Theorem 7.0.0.2], see below for an explanation of this analogy. We also deduce the following dévissage theorem for which is a special case of Theorem 6.1 (ii).
Theorem 0.2.
Let be an exact fully faithful functor between small abelian categories such that generates via extensions (equivalently, satisfies the assumptions of Quillen’s Dévissage theorem). Then induces an equivalence of spectra
The proof of Theorem 0.1 is based on the following much stronger version of Barwick’s theorem of the heart [Bar15, Theorem 6.1]. This is a special case of Theorem 4.1.
Theorem 0.3.
Let and be small stable categories with bounded -structures resp. Let be an integer and suppose that we have an exact and -exact functor such that the following conditions hold:
the image generates as an idempotent-complete stable subcategory;
for the map is an isomorphism for and a monomorphism for
Then the induced map is an isomorphism for and a monomorphism for
We mention the following immediate corollaries, which are again special cases of more general results for large categories (Corollary 3.2 and Theorem 6.1 (i)).
Corollary 0.4.
Let be a small stable category with a bounded -structure Suppose that for some the maps are isomorphisms for and (this always holds for ). Then the map is an isomorphism for and a monomorphism for
Note that in the situation of the above corollary the maps are automatically monomorphisms for
Corollary 0.5.
Let be a fully faithful exact functor between small abelian categories, such that generates via extensions. Then the induced map is an isomorphism for and a monomorphism for
We note that the above results are new only when the hearts are non-noetherian (and non-artinian). Namely, if is noetherian, then we have by [AGH19, Theorem 3.6], and we have (more precisely, for all ).
We prove that the estimates in Corollaries 0.4 and 0.5 are sharp (Theorem 7.1), even if we are dealing with small dg categories resp. small abelian categories over a field. In particular, we obtain that the naive -theoretic theorem of the heart fails for which is a much more precise version of a result of Ramzi-Sosnilo-Winges [RSW24, Theorem B].
We now explain a very surprising analogy with the familiar results on connective -ring spectra. Intuitively the notion of a -structure is dual to the notion of a weight structure [Bon10, Sos19]. In particular, the bounded -structures are “dual” to bounded weight structures. Consider an idempotent-complete stable category with a bounded weight structure For simplicity we assume that is generated by a single object (this assumption is harmless, it will make no essential difference in what follows). In this case we may assume that is of weight zero. Putting (meaning the spectrum of endomorphisms), we obtain and is a connective -ring. Conversely, for a connective -ring the category has a unique bounded weight structure such that has weight zero [Sos19].
The following statement is well-known. It is due to Waldhausen [Wal78] for simplicial rings and connective -theory, and the general case is proven by Land and Tamme [LT19] (the case follows from the proof of [LT19, Corollary 3.5]).
Proposition 0.6.
It is clear from the formulations that our Theorem 0.3 for some is “dual” to Proposition 0.6 for both in the assumptions and in the conclusion one simply replaces with for the relevant spectra, and one replaces epimorphisms with monomorphisms. Continuing this analogy, Corollary 0.4 for some is “dual” to the special case of Proposition 0.6 for the same when and for Further, Corollary 0.5 is “dual” to the case and We do not have any conceptual explanation for why such a duality takes place, but we note that in both contexts the (co)connectivity estimates are sharp.
Now we explain why from this point of view our Theorem 0.1 is “dual” to the Dundas-Goodwillie-McCarthy theorem. We recall some notation. We denote by the category of idempotent-complete stable categories and consider the universal finitary localizing invariant [BGT13, E25c]. The category is naturally symmetric monoidal, and moreover it is rigid (in the sense of Gaitsgory and Rozenblyum [GaiRoz17, Definition 9.12]) by [E25c, Theorem 0.1]. We also consider the universal finitary -invariant localizing invariant so is a quotient of by the cocomplete ideal generated by the reduced motive of the flat affine line By [E25c, Theorem 8.1] this is a smashing localization. More precisely we have the idempotent -algebra in given by the geometric realization where We have an equivalence
The inclusion functor has a left adjoint and a right adjoint given by
where is the internal in We also denote by the functor of global sections, i.e. where is the unit object
By [E25c, Theorem 9.3] (whose proof of course uses the DGM theorem) for a connective -ring we have a natural equivalence of spectra
where and is the topological cyclic homology [BHM93, NS18]. The map is the cyclotomic trace. In view of the isomorphism (0.2), the DGM theorem states that we have an equivalence
More precisely, [DGM13, Theorem 7.0.0.2] states a more general version for a quotient of by a nilpotent ideal, but this is automatic by [LT19, Corollary 3.5].
Now, by [E25c, Corollary 8.3] for we have
Therefore, the statement of Theorem 0.1 is “dual” to the DGM theorem if we continue the above analogy. This was in fact our motivation for why theorem of the heart should hold for
The original goal of this paper was much more modest: to prove the generalizations of Barwick’s theorem of the heart [Bar15, Theorem 6.1] and Quillen’s Dévissage theorem [Qui73, Theorem 4] for large categories. However, the methods which we developed allowed for much better results. We briefly explain which large abelian categories we consider, and which -structures on large categories are “nice” in this context.
A Grothendieck abelian category is coherently assembled if the colimit functor has an exact left adjoint Recall that the mere existence of this left adjoint means that the category is compactly assembled [Lur18, Definition 21.1.2.1] (since is an ordinary category, this means that is continuous in the sense of Joyal and Johnstone [JJ82]). In our situation this is equivalent to the condition that satisfies the Grothendieck’s axiom (AB6) [Gro57]. As explained in [E24, Proposition E.8], if is a compactly generated Grothendieck abelian category, then is coherently assembled if and only if is locally coherent (i.e. equivalent to an ind-completion of a small abelian category).
Recall that if is a Grothendieck abelian category, then its unseparated derived category is identified with the (dg nerve of) the dg category of unbounded complexes of injective objects of see [Kr05, Kr15] and [Lur18, Section C.5.8]. If is moreover coherently assembled, then the category is dualizable (Propositions 1.26 and 1.34). Hence, we can define the continuous -theory of
If is locally coherent, then by [Kr15, Theorem 4.9], hence
Now, if is a dualizable (presentable stable) category, then a -structure is called compactly assembled if it is accessible, compatible with filtered colimits and the functor (left adjoint to ) is -exact. In this case the heart is automatically coherently assembled. We further impose a condition that the -structure is continuously bounded, which means that is generated by (equivalently, by ) as a localizing subcategory. For brevity we say that is a dualizable -category if it is a dualizable category equipped with a compactly assembled continuously bounded -structure.
For a dualizable -category we have a (strongly continuous) realization functor which follows from the universal property of explained in [Lur17b], and also from the universal property of from [Lur18, Proposition C.5.5.20, Theorem C.5.8.8]. If is compactly generated, then the -structure on induces a bounded -structure on and the functor is identified with the ind-completion of the realization functor With this in mind, it is clear that the above statements for small stable categories with bounded -structures can at least be formulated more generally for dualizable -categories.
However, one has to be cautious with generalizations to the dualizable context. Namely, for a small stable category with a bounded -structure we have by a theorem of Antieau-Gepner-Heller [AGH19, Theorem 2.35]. However, for a dualizable -category in general we have even if the realization functor is an equivalence (Corollary 8.2). This is a very rare example of a situation when a reasonable statement about dualizable categories fails in general, but holds for compactly generated categories.
We mention the following natural questions which are not studied in this paper. The first one is somewhat vague, and it is motivated by the paper [LS25] and by the above analogy (duality) with weight structures.
Question 0.7.
Let be a dualizable category with a -structure satisfying the following conditions: the -structure is accessible, compatible with filtered colimits, is generated by as a localizing subcategory, and for some the functor is left -exact. The (dualizable versions of) the above results deal with the case Is there a reasonable generalization to the case
The following question is motivated by the above analogy with connective -rings. We introduce some non-standard notation. Let be a (not necessarily exact) functor between -presentable stable categories, commuting with -filtered colimits. For denote by the universal polynomial functor of degree commuting with -filtered colimits, with a map (its existence follows from the adjoint functor theorem). We say that the coanalytic approximation of is the colimit (computed objectwise).
Question 0.8.
Let be a dualizable category (without a -structure), and let be a colimit-preserving endofunctor. We consider the tensor algebra as a monad on and take the category of modules which is dualizable by [E24, Proposition C.1]. Consider the functor
We denote by
the coanalytic approximation of (0.3).
Now suppose that is a dualizable -category and is (colimit-preserving and) left -exact. Is it true that the map
is an equivalence of spectra?
We give a brief overview of the structure of the paper.
In Section 1 we introduce and study locally -coherent exact -categories, focusing on coherently assembled exact -categories. We are mostly interested in the presentable stable envelope of a coherently assembled exact -category and in the properties of this category, such as dualizability. In particular, we define the continuous -theory of such exact -categories by putting Our approach in this section is somewhat minimalistic, in particular we mostly avoid the non-coherently assembled case. We reduce most of the non-trivial statements to the familiar results about small exact -categories and their (small) stable envelopes. Among other things, we explain a natural version of the situation of dévissage for coherently assembled abelian categories in Subsection 1.9. The results of this section are used in the proofs of our main theorems, especially in Section 3.
In Section 2 we introduce and study compactly assembled -structures on dualizable categories, mostly focusing on continuously bounded -structures. In particular, in Subsection 2.4 we explain a general method of constructing a nice -structure via a functor from a coherently assembled abelian category, based on a similar result for small categories. This method allows to construct many auxiliary categories with -structures, which are used essentially in all the proofs of the main theorems. In particular, we consider the presentable stable envelopes with natural -structures with the same heart. Here is considered as an exact -category with the induced exact structure from Again, our approach in this section is minimalistic, and some statements can be generalized to the case when is merely presentable stable, the -structure is compatible with filtered colimits and right complete, and the Grothendieck prestable category is anticomplete in the sense of [Lur18, Definition C.5.5.4].
In Section 3 we prove refined theorem of the heart in the context of dualizable categories (Theorem 3.1). It states that for a dualizable -category and for the spectrum is -coconnective. This in particular implies the above Corollary 0.4. One of the key ingredients is the surjectivity of the map which is proved by a very non-trivial argument. In particular, it requires some careful analysis of the (compactly generated) iterated Calkin categories and natural -structures on them. Another important ingredient is the construction of certain short exact sequences of dualizable categories (Proposition 3.7), which relate the categories for various
In Section 4 we prove Theorem 4.1 on the coconnectivity estimates, which is a generalization of Theorem 0.3. One of the main ideas is a very abstract construction: given a strongly continuous exact functor between dualizable categories such that generates there is a canonical sequence in such that and each monad on is obtained as a so-called deformed tensor algebra. If and are dualizable -categories and is -exact, then all are also naturally dualizable -categories and the transition functors are -exact. A further analysis of deformed tensor algebras allows to eventually reduce Theorem 4.1 to Theorem 3.1 (refined theorem of the heart). We also deduce similar results for higher nil groups (Corollaries 4.11 and 4.12).
In Section 5 we prove theorem of the heart for for dualizable -categories (Theorem 5.1), which is a generalization of Theorem 0.1. Here we briefly explain a more precise statement, for simplicity restricting to the case of small categories. For denote by the -th term of the standard filtration on namely is obtained by taking the colimit over in the right hand side of (0.1). Then the localizing invariants form a direct sequence with and We prove by induction on that if satisfies the assumptions of Theorem 0.3 for some then the map of spectra
is null-homotopic for
The case directly implies Theorem 0.1.
In Section 6 we deduce dévissage theorems for and of coherently assembled abelian categories, which are generalizations of Theorem 0.1 and Corollary 0.5.
In Section 7 we prove that our estimates in Corollaries 0.4 and 0.5 are sharp (Theorem 7.1), working over a field The constructions are based on the exact category of vector bundles on a (projective) cuspidal cubic curve over and on the abelian category of effaceable functors Essentially we only use that is proper and We also prove the sharpness of the estimates for higher nil groups (Theorem 7.3), also giving a non-trivial example of a computation of for an abelian category
In Section 8 we study examples of compactly assembled -structures and coherently assembled abelian categories. In particular, we consider sheaves on locally compact Hausdorff spaces (Proposition 8.1). As a special case, we obtain that the abelian category of sheaves of vector spaces on the real line is coherently assembled, and we have Therefore, already the Schlichting’s vanishing theorem for of small abelian categories [Schl06, Theorem 6] does not generalize to coherently assembled abelian categories. We consider the category of nuclear solid modules over (the original version due to Clausen and Scholze [CS20]), and show that it has a natural compactly assembled continuously bounded -structure (Proposition 8.3). We also prove a generalization of Chase criterion [Ch60] of coherence of associative unital rings. This is Theorem 8.5. One way to state it is as follows: a dualizable -module in is a coherently assembled abelian category if and only if the category of flat objects in has infinite products. Here an object is flat if the functor is left exact. We recall that an -module in is the same thing as a presentable additive (ordinary) category By the forthcoming work of Levy, Liang and Sosnilo [LLS], is dualizable if and only if is a Grothendieck abelian category satisfying (AB6) and (AB4*). Equivalently by Roos’ theorem [Roo65] this means that is equivalent to a category of almost modules, i.e. where is an associative unital ring and is a two-sided ideal such that Finally, in Subsection 8.4 we mention some example and questions which are not covered by this paper.
Convention. We will freely use the theory of -categories as developed in [Lur09, Lur17a]. We will simply say “category” instead of -category. If the category is -truncated (i.e. the morphism spaces are discrete), we will say “ordinary category”. Since a stable category is naturally enriched over spectra, for we denote by the spectrum of morphisms. We denote by the category of spectra, and by the category of connective spectra. As we already mentioned above, we denote by the category of idempotent-complete small stable categories and exact functors. Throughout this paper we use the homological grading, unless otherwise stated.
Acknowledgements. I am grateful to Dmitry Kaledin, Ishan Levy, Maxime Ramzi and Vladimir Sosnilo for useful discussions. This work was performed at the Steklov International Mathematical Center and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2025-303). I was partially supported by the HSE University Basic Research Program.
1. Coherently assembled abelian categories and exact categories
1.1. Reminder on exact categories
Definition 1.1.
[Bar15, Section 3] An exact category is an additive category with a choice of two subcategories consisting of (exact) inclusions resp. (exact) projections, such that the following conditions hold.
For the map is an inclusion and the map is a projection.
The class of inclusions is stable under pushouts along all maps, and the class of projections is stable under pullbacks along all maps.
For a square
in the following are equivalent:
(a) is an inclusion, is a projection and the square is cocartesian;
(b) is an inclusion, is a projection and the square is cartesian.
Such a square is called an exact square. If we will say that is a short exact sequence in
We will omit the subcategories and from the notation, assuming that they are chosen. An additive functor between exact categories is called exact if it preserves short exact sequences, or equivalently exact squares. We denote by the (additive) category of exact functors We denote by the category of small exact categories and exact functors.
We say that an exact functor between exact categories reflects exactness if for any square in if its image is exact in then it is exact in
If we denote by the category of small additive categories and additive functors, then the forgetful functor has a left adjoint Here is the exact category where consists of split monomorphisms and consists of split epimorphisms. We call this a split exact structure on Note that if is a small exact category, then the identity functor is exact, but it does not reflect exactness in general.
Note that any stable category can be considered as an exact category with For stable categories exactness of a functor in the above sense is equivalent to the exactness in the usual sense. Denoting by the category of small stable categories and exact functors, we obtain a fully faithful functor By [Kle22, Proposition 4.22, Corollary 4.24] it has a left adjoint, which we denote by The category is called a stable envelope of the category The following is due to Klemenc.
Proposition 1.2.
[Kle22, Propositions 4.17, Proposition 4.25] Let be a small exact category, and denote by the universal exact functor. Then is fully faithful and reflects exactness, and its essential image is closed under extensions.
We refer to [SW26] for further details. We will need the following more precise statement about the stable envelope of an exact category First, consider the stable category with the universal additive functor We have
where the superscript “add” means that we consider the additive functors. We identify with the full subcategory of this category of functors, and we have Clearly, we have a quotient functor Its kernel is denoted by and it is generated as a stable subcategory by the objects of the form where is a short exact sequence in and
As explained for example in [SW26, Proposition 3.9], the associated functor takes values in and we have
Such functors are called effaceable, and they form an ordinary additive category, denoted by This category is in fact abelian, more precisely the following holds.
Proposition 1.3.
Remark 1.4.
If is an ordinary exact category, then the above category is denoted by in [Nee21], since it is the category of bounded acyclic complexes. This should not lead to confusion, because we will not consider the unbounded version.
We also recall a more general class of functors. Recall that for an exact category an additive functor is called weakly effaceable if for any and for any there exists a projection in such that We denote by the category of weakly effaceable functors It is easy to see that we have Moreover, by [SW26, Proposition 3.9] an additive functor is effaceable if and only if it is finitely presented and weakly effaceable. It is easy to see (assuming Proposition 1.3) that the category is abelian and locally coherent, and its full subcategory of finitely presented objects coincides with Indeed, it suffices to observe that any map from a finitely presented additive functor to a weakly effaceable functor factors through an effaceable functor.
We will use the following notation. If is a stable category, then for and we put For a small exact category and for we put where is the universal exact functor. By Proposition 1.2 we have for Also, the abelian group classifies extensions of by in
The following is a well-known observation, but we are not aware of a precise reference.
Proposition 1.5.
Let be a stable category. Let be a small full additive subcategory closed under extensions. We consider as an exact category with the induced exact structure from Consider the exact functor induced by the inclusion Then for the induced map is an isomorphism for and a monomorphism for
Proof.
The isomorphisms for follow from Proposition 1.2 and from the description of via extensions. Now take an element By [SW26, Lemma 4.3] there exists a projection such that we have Putting we see that where If the image of in is zero, then the map in factors through hence is contained in the image of the map But this exactly means that This shows the injectivity of the map ∎
We recall the following class of exact subcategories. The condition in the following definition already appears in [Kel96].
Definition 1.6.
[SW26, Definition 1.1] Let be an exact category. A full additive subcategory is called left special if is closed under extensions in and for any projection in with there exists and a map such that the composition is a projection in
Theorem 1.7.
[SW26, Theorem 4.5] Let be a small exact category, and let be a left special full subcategory. We consider as an exact category with an induced exact structure. Then the functor is fully faithful.
For completeness we mention the following statement, which might be not in the literature. In the case of ordinary exact categories this is [BS01, Theorem 2.8], with a different proof.
Proposition 1.8.
Let be an idempotent-complete exact category. Then the category is also idempotent-complete.
Proof.
Denote by the idempotent completion. The category is idempotent-complete by [Bon10, Theorem 5.3.1, Proposition 6.2.1], so we have a short exact sequence in
By Proposition 1.3 the category has a bounded -structure, hence by [AGH19, Theorem 2.35] we have The long exact sequence gives the surjectivity of the map By [Th97, Theorem 2.1] this implies that is idempotent-complete. ∎
1.2. Accessibly exact categories
To deal with large exact categories we will need the following basic construction. We will use the following notation: if is a small category and is a regular cardinal, then we write the objects of as where is a -directed poset and we are assuming a functor One can more generally take to be a -filtered -category, but recall that by [Lur09, Proposition 5.3.1.18] for such there exists a -directed poset and a cofinal functor
Proposition 1.9.
Let be a regular cardinals. Let be a small exact category. Then the category has the following exact structure. A morphism in is an inclusion resp. a projection if it is a -small -directed colimit of inclusions resp. projections in
Taking the union over we obtain an exact structure on The (fully faithful) functor preserves and reflects exactness. Its essential image is closed under extensions.
Proof.
Consider the stable envelope and denote by the universal exact functor. By Proposition 1.2 is fully faithful, reflects exactness, and its essential image is closed under extensions in Consider the stable category Then we have a fully faithful functor
We claim that the essential image of is closed under extensions. Indeed, consider a cofiber sequence of the form in By [Lur09, Proposition 5.3.5.15] and its proof the morphism is a -small -directed colimit of morphisms where Hence, is contained in the essential image of since each object is contained in the essential image of This shows that the essential image of is closed under extensions.
Consider the induced exact structure on The same argument as above shows that short exact sequences in this category are exactly the -small -directed colimits of short exact sequences in
Finally, the assertions about the functor follow from the above. ∎
Definition 1.10.
We define a locally -coherent exact category to be a category of the form where is a small exact category and is a regular cardinal. The exact structure is described in Proposition 1.9.
An accessibly exact category is a locally -coherent exact category for some
We will consider the accessible stable categories as accessibly exact categories in the natural way.
Proposition 1.11.
Let be a small exact category and let be a regular cardinal.
The inclusion functor preserves and reflects exactness. Its essential image is closed under extensions.
Let be an additive subcategory closed under extensions, equipped with the induced exact structure from Then the functor preserves and reflects exactness. Its essential image is closed under extensions.
Proof.
(i) If is stable with the standard exact structure, then the statement is evident. The general case reduces to the stable case using Proposition 1.9. Namely, we have a commutative square
and both vertical functors are fully faithful, preserve and reflect exactness, and their essential images are closed under extensions. The same holds for the lower horizontal functor by the above special case. Hence, the same holds for the upper horizontal functor.
(ii) The proof is similar, but a little caution is required: the functor does not have to be fully faithful. Instead, we first observe that we can replace by by Proposition 1.9. Assuming that is stable, the same argument as in loc. cit. shows that the essential image of in is closed under extensions. It also shows that this functor reflects exactness. ∎
The following definition is essentially tautological, but we spell it out for completeness.
Definition 1.12.
Let be an accessibly exact category and let be a regular cardinal such that has -filtered colimits. We say that the -filtered colimits are exact in if the class of short exact sequences in is closed under -filtered colimits.
This exactness holds in the following basic situation.
Proposition 1.13.
Let be a locally -coherent exact category. Then the -filtered colimits are exact in
Proof.
Again, the stable case is evident. The general case reduces to the stable case by Proposition 1.9: the functor commutes with -filtered colimits and reflects exactness. ∎
The following observation is slightly subtle.
Proposition 1.14.
Let be an accessibly exact category and let be a regular cardinal. Suppose that has exact -filtered colimits. Then the full subcategory is closed under extensions.
Proof.
Let be a short exact sequence in and suppose that Let be a -directed system in We need to prove that the map is an equivalence of spaces. By assumption this holds if we replace with or Using the long exact sequences and the five-lemma, we reduce the question to showing that the map of abelian groups
is a monomorphism.
Take some and an element It corresponds to a short exact sequence For consider the pushout Since -filtered colimits are exact by assumption, the image of in corresponds to the short exact sequence Suppose that this extension splits, i.e. we have a section Then factors through for some This means that the image of in vanishes. This proves the injectivity of (1.2) and the proposition. ∎
In the situation of Proposition 1.14 we consider as an exact category with the induced exact structure from We make further observations about the subcategories for various
Proposition 1.15.
Let be a locally -coherent exact category and let be a regular cardinal.
For any the functor commutes with -filtered colimits.
The inclusion functor is left special. In particular, the functor is fully faithful.
Proof.
We prove (i). If is -accessible stable with standard exact structure, then the statement is evident. The general case reduces to the stable case by the proof of Proposition 1.9: we have a fully faithful -accessible functor and This proves (i).
The second assertion of (ii) follows from the first one by Theorem 1.7. We prove the left speciality. Consider a short exact sequence in with Denote by the corresponding element. Let where is -small and -directed, and for By (i) there exists some and an element which maps to We obtain the associated short exact sequence By Proposition 1.14 we have By construction, the projection factors through which proves that is left special in ∎
In the situation of Proposition 1.15 we will consider as a directed union of over all regular
1.3. Grothendieck exact categories
Definition 1.16.
Let be an accessibly exact category. We say that is a Grothendieck exact category if has exact filtered colimits.
It is convenient to use the following convention: if is a locally -coherent exact category, then is a union of exact categories for all regular
Proposition 1.17.
Let be a locally -coherent exact category with filtered colimits. Suppose that the colimit functor is exact. Then is a Grothendieck exact category.
Proof.
We need to show that the functor is exact. By Proposition 1.13 for any small exact category the colimit functor is exact. Hence, so is the composition
This proves the proposition. ∎
Example 1.18.
A Grothendieck abelian category can be considered as a Grothendieck exact category.
Example 1.19.
Let be a small exact category. Then is a Grothendieck exact category by Proposition 1.13.
Example 1.20.
Any presentable stable category can be considered as a Grothendieck exact category. More generally, any Grothendieck prestable category [Lur18, Definition C.1.4.2] is a Grothendieck exact category.
1.4. Criterion of -accessibility for exact categories
We will use the following general characterization of locally -coherent exact categories.
Proposition 1.21.
Let be an accessibly exact category and let be a regular cardinal. Then is locally -coherent if and only if the following conditions hold.
is -accessible (as an abstract category).
The -filtered colimits are exact in
For the functor commutes with -filtered colimits.
Moreover, in this case the full subcategory is closed under fibers of projections in
Proof.
We first prove the “only if” direction. (i) holds by definition, (ii) holds by Proposition 1.13 and (iii) holds by Proposition 1.15. We prove the “moreover” assertion. Let be a short exact sequence in with Consider the functor from Proposition 1.9. Then the objects are -compact, hence so is Since is fully faithful and commutes with -filtered colimits, it follows that is -compact in
Now we prove the “if” direction. By (ii) and by Proposition 1.14 the full subcategory is closed under extensions. We consider as a small exact category with the induced exact structure from We equip with the exact structure from Proposition 1.9. By (ii) the natural functor is exact. By (i) is an equivalence (of abstract categories). It remains to prove that reflects exactness. In other words, we need to prove that the class of short exact sequences in is the smallest class which contains short exact sequences in and is closed under -filtered colimits.
Let be a short exact sequence in and denote by the corresponding element. Let where is -directed and Putting we see that is a -directed colimit of short exact sequences Hence, we may and will assume that Let where is -directed and By (iii) there exists and an element which maps to Consider the corresponding short exact sequence For put By Proposition 1.14 we have We conclude that is the -directed colimit of over which proves that reflects exactness. ∎
1.5. Presentable stable envelopes
Let be a Grothendieck exact category, and let is presentable stable category. Denote by the category of continuous (i.e. filtered colimit-preserving) exact functors
Definition 1.22.
For a Grothendieck exact category we denote by a presentable stable category with a continuous exact functor such that for any the precomposition with induces an equivalence
We call the presentable stable envelope of
We use the notation because this is a generalization of the unseparated derived category of a Grothendieck abelian category see Proposition 1.26 below.
Proposition 1.23.
For a Grothendieck exact category its presentable stable envelope exists and it is given by
Proof.
Let us temporarily denote by the right hand side of (1.3). Choose a regular cardinal such that is locally -coherent exact. Then restriction to defines a fully faithful functor
Its essential image consists of functors which commute with -small cofiltered colimits. It follows that is presentable stable, and is accessible and commutes with infinite products. Therefore, has a left adjoint Consider the composition
Then the kernal of is generated (as a localizing subcategory) by objects of the form where is a -small directed poset and , is a functor.
Now define to be the left Kan extension of Then is exact. We further define
We claim that satisfies the required universal property. First, note that commutes with filtered colimits. To check this, we only need to see that commutes with -small directed colimits, and this follows from the above description of the kernel of
Now for any the functor induces an equivalence
where the superscript “” means the commutation with -filtered colimits. It follows formally that induces an equivalence
The case of locally coherent exact categories is studied in [NW25].
Proposition 1.24.
Let be a locally coherent exact category. Then we have an equivalence
Proof.
The description of follows from the universal property, see also [NW25, Corollary 2.24]. ∎
The following basic example is familiar from [Lur18].
Proposition 1.25.
Let be a Grothendieck prestable category, which we consider as a Grothendieck exact category. Then we have an equivalence compatible with the universal functors from
Proof.
This follows from the universal properties. Namely, for a presentable stable category a functor is continuous and exact if and only if commutes with colimits. ∎
Recall that for a Grothendeick abelian category the so-called unseparated derived category can be described as a (dg nerve of) the dg category of complexes of injective objects of It has a natural right complete -structure, compatible with filtered colimits, and we have Moreover, is generated by the heart as a localizing subcategory. We refer to [Kr05, Kr15] and [Lur18, Section C.5.8] for a detailed account.
Proposition 1.26.
For a Grothendieck abelian category we have
Proof.
This is explained in [Lur17b]. ∎
Remark 1.27.
Essentially the same argument as in loc. cit. proves a similar result for the so-called Grothendieck abelian -categories, i.e. categories of the form where is a prestable Grothendieck category and [Lur18, Definition C.5.4.1]. Namely, if we equip with the induced exact structure from then is equivalent to where is left adjoint to Here resp. is the category of Grothendieck prestable categories resp. Grothendieck abelian -categories, where the -morphisms are colimit-preserving left exact functors.
We will crucially use such categories below, but only in the relatively elementary case when is the connective part of a compactly assembled -structure on a dualizable category (Definition 2.8). We will not need the above description, instead we will apply the general machinery of coherently assembled exact categories from Subsection 1.7.
1.6. Reminder on compactly assembled and dualizable categories
We recall some basic notions from [Lur18] and [E24], and also fix the notation and terminology. Recall that we simply say “category” instead of “-category”, and say “ordinary category” when the spaces of morphisms are discrete.
If is an accessible category with filtered colimits, then is called compactly assembled if the colimit functor has a left adjoint This is [Lur18, Definition 21.1.2.1]. By [E24, Proposition 1.24] any compactly assembled category is -accessible, hence the functor takes values in (the latter follows for example by the same argument as in [E24, Corollary 1.22]).
If is a compactly assembled category, then a morphism in is called compact if the map in factors through where is the Yoneda embedding. Recall that any compact morphism in is (homotopic to) a composition of two compact morphisms, by the same argument as in [E24, Corollary 1.41].
A functor between compactly assembled categories is called continuous if it commutes with filtered colimits. Further, is called strongly continuous if it is continuous and the following (lax commutative) square commutes:
Note that if has a right adjoint then the strong continuity of exactly means that is continuous (the same proof as in [E24, Proposition 1.14]). In general, the strong continuity of a continuous functor exactly means that preserves compact morphisms (it suffices to check this for morphisms between the objects of ).
We recall that denotes the category of presentable categories and colimit-preserving functors. It is symmetric monoidal with the Lurie tensor product, and the unit object is given by the category of spaces We denote by the full subcategory of presentable stable categories, which is also the category of modules over the idempotent -algebra In particular, is closed under tensor products in and the unit object is given by A presentable stable category is called dualizable if it is a dualizable object in We will simply say “dualizable category”, assuming that it is presentable and stable.
By [Lur18, Proposition D.7.3.1], is dualizable if and only if it is compactly assembled. In particular, the above notions apply. In this paper we will deal with some functors between dualizable categories, which are strongly continuous but not exact, such as a truncation endofunctor for a nice -structure. For this reason we will specify if the functor is supposed to be exact. We denote by the non-full subcategory of dualizable categories and strongly continuous exact functors. For we denote by the (idempotent-complete small stable) category of strongly continuous exact functors.
We recall that again by [Lur18, Proposition D.7.3.1] a presentable stable category is dualizable if and only if it is a retract in of a compactly generated category. We have a fully faithful embedding
We will use the convention from [E25c] for Calkin categories. Namely, for we put
This is a compactly generated (presentable stable) category. We have
where the category is introduced in [E24, Definition 1.60]. In this paper we will simply write for the latter category.
Finally, we recall the continuous -theory and more general (accessible) localizing invariants for dualizable categories, as defined in [E24]. In this paper we consider the -valued localizing invariants [BGT13]. For such and for we put
Then is a localizing invariant by [E24, Proposition 4.6] and we have a functorial isomorphism for by [E24, Proposition 4.7]. Moreover, is uniquely determined by these properties by [E24, Theorem 4.10].
1.7. Coherently assembled exact categories
The following definition specifies the class of large exact categories for which one can define -theory and other localizing invariants.
Definition 1.28.
Let be a Grothendieck exact category. We say that is coherently assembled if is compactly assembled and the functor is exact.
We denote by the category of coherently assembled exact categories, where the -morphisms are strongly continuous exact functors.
Remark 1.29.
Example 1.30.
Let be a dualizable (presentable stable) category with the standard exact structure. Then is coherently assembled. This gives a fully faithful functor
We will see below that a coherently assembled exact category is always locally -coherent. First we observe an equivalent characterization.
Proposition 1.31.
Let be a Grothendieck exact category. The following are equivalent.
is coherently assembled.
There exists a locally coherent exact category and a retraction such that both and are exact and commute with filtered colimits.
In particular, if is locally coherent, then it is coherently assembled.
Proof.
Note that the final assertion is obvious: if is locally coherent, then is obtained by applying to the exact inclusion
Suppose that is locally -coherent. Consider the retraction Both functors commute with filtered colimits. By Remark 1.29 the functor is exact, and by definition of a Grothendieck exact category the functor is exact. This proves the implication.
Since and are continuous, the functor is isomorphic to the composition
Here the first and the third functors are exact by assumption, and the second functor is exact by the above observation on the locally coherent case. This proves the proposition. ∎
We explain a natural analogue of split exact structures in the context of large exact categories, see also Remark 1.33 below.
Proposition 1.32.
Let be a compactly assembled additive category. Then has an exact structure such that a map in is an exact projection if and only if any compact morphism factors through With this exact structure is a coherently assembled exact category. Moreover, short exact sequences in are exactly the filtered colimits of split short exact sequences.
Proof.
We first consider the case when is compactly generated, and put Then is identified with the category of flat additive functors this is a straightforward generalization of [Lur17a, Theorem 7.2.2.15]. The definition of an exact projection in this case gives the class of morphisms which are effective epimorphisms in the ambient category Hence, we simply obtain the standard exact structure on
Now consider the general case. Choose a retraction where is additive and compactly generated, and are additive and continuous, and is moreover strongly continuous (e.g. take ). Denote by the category of cofiltered limit-preserving functors As in Proposition 1.23 we see that we have a universal continuous additive functor Consider also a similar functor The latter is simply the ind-completion of the functor where is equipped with the split exact structure. It follows from the retraction that is fully faithful and its essential image is closed under extensions. Consider the induced exact structure on Using the retraction again we see that is a coherently assembled exact category, and any short exact sequence in is a filtered colimit of split short exact sequences. It remains to obtain the stated description of exact projections in
If is an exact projection in then is a filtered colimit of split exact projections Any compact morphism factors through some hence through hence through
Conversely, left be a morphism in such that any compact morphism factors through Let where is directed. Each map is compact, hence so is by the strong continuity of Moreover, for each the map factors through It follows from the above special case that is an exact projection in and applying we conclude that is an exact projection in ∎
Remark 1.33.
The category from the proof of Proposition 1.32 has a natural -structure such that the connective part is identified with the category of cofiltered limit-preserving functors Moreover, is a dualizable module over i.e. a dualizable object of – the symmetric monoidal category of additive presentable -categories [LLS]. Moreover, any dualizable additive category can be obtained in this way: one recovers as a full subcategory of flat objects of Here an object is flat if the functor commutes with finite limits.
Next we show that the presentable stable envelope is well-behaved.
Proposition 1.34.
Let be a coherently assembled exact category. The category is dualizable and the functor is strongly continuous and fully faithful. Moreover, the essential image of in is closed under extensions, and reflects exactness.
Let be a strongly continuous exact functor between coherently assembled exact categories. Then the induced functor is also strongly continuous. Therefore, we have a well-defined functor It is left adjoint to the inclusion.
Proof.
(i) We first consider the case when is locally coherent. Then by Proposition 1.24 the category is compactly generated, hence dualizable. The strong continuity is immediate: the functor is identified with The fully faithfulness follows from Proposition 1.2. The remaining assertions follow from Proposition 1.9.
Now consider the general case. By Proposition 1.31 we have a retraction where is locally coherent and both and are exact and continuous. Now, both functors and are continuous, and the functor is a retract of the functor Hence, all the assertions about and follow from the above special case when is locally coherent.
(ii) Put We need to show that for we have an isomorphism It suffices to consider the case where and is the universal exact functor. Then the isomorphism follows directly from (i): we have a commutative square of compactly assembled categories
in which both vertical arrows and the upper horizontal arrow are strongly continuous.
It follows formally that the functor is left adjoint to the inclusion: the adjunction counit is well-defined by (i). ∎
The above proposition allows to define the continuous -theory and other localizing invariants (such as ) for coherently assembled exact categories.
Definition 1.35.
Let be a localizing invariant. For a coherently assembled exact category we put
This defines a functor
Next, we show the expected local -coherence.
Proposition 1.36.
Let be a coherently assembled exact category. Then is locally -coherent.
Proof.
It suffices to check that the conditions from Proposition 1.21 are satisfied. Since is compactly assembled, it is -accessible by [E24, Proposition 1.24]. Next, all filtered colimits are exact in in particular, -filtered colimits are exact. It remains to prove that for the functor commutes with -filtered colimits. By Proposition 1.34 the functor is strongly continuous, in particular, it preserves -compact objects. Hence, we have It follows that the functor commutes with -filtered colimits by stability. Applying Proposition 1.34 again, we see that for we have an isomorphism Hence, the functor also commutes with -filtered colimits. ∎
From now on, for a coherently assembled exact category we consider the small category as an exact category with the induced exact structure from and similarly for any regular Note that by Proposition 1.15 for any uncountable regular cardinals the functor is fully faithful, and we consider as a directed union of
Proposition 1.37.
Let be a coherently assembled exact category. Then the natural functor is fully faithful.
Proof.
By Proposition 1.31 we have a retraction where is locally coherent and both and are continuous and exact. Hence, the functor is a retract of the functor Thus, we may and will assume that is locally coherent.
Consider the functor which is fully faithful by Proposition 1.34. Let be an uncountable regular cardinal. By Proposition 1.15 it suffices to prove that the functor is fully faithful. Take the restriction By [SW26, Lemmas 4.2 and 4.3] it suffices to prove that for and for the functor is weakly effaceable. Namely, we need to show that for any and for any element there exists a projection in such that the image of in is zero.
We first consider the case Let where is directed and By Proposition 1.24 is compact in hence there exists such that is the image of some element By [SW26, Lemma 4.3] there exists a projection in such that the image of in vanishes. Then the image of in vanishes, as required.
Next, if is a -small direct sum of some then it suffices to take to be the direct sum of where the projection is as above. Now, consider the general case, i.e. where is a -small directed poset, and It suffices to prove that the map
is a projection in By Proposition 1.21 it suffices to prove that (1.4) is a projection in We may and will assume that the exact structure on is split, i.e. we have only split short exact sequences. Then is identified with the category of flat additive functors with the induced exact structure from the Grothendieck prestable category The class of flat objects in is closed under fibers of effective epimorphisms, which implies that (1.4) is a projection in as required. ∎
The following corollary is immediate.
Corollary 1.38.
Let be a coherently assembled exact category, and let be a presentable stable category. Suppose that we have a continuous exact functor which is fully faithful, reflects exactness and its essential image is closed under extensions. Consider the induced (continuous exact) functor and denote by the universal continuous exact functor, so that . Then for the map
is an isomorphism for and a monomorphism for
Proof.
Finally, we prove a version of [SW26, Theorem 4.5], which is deduced from loc. cit.
Proposition 1.39.
Let be a functor between coherently assembled exact categories, which is fully faithful, strongly continuous, reflects exactness, and such that the essential image of is closed under extensions. Suppose that the inclusion is left special (Definition 1.6). Then the induced functor is fully faithful.
Proof.
The functor is a retract (in ) of the functor which is fully faithful by Theorem 1.7. ∎
Remark 1.40.
In Proposition1.39 one can equivalently assume that is left special in However, we prefer to formulate the condition in terms of small categories.
1.8. Coherently assembled abelian categories
Definition 1.41.
A Grothendieck abelian category is coherently assembled if it is such as an exact category. This means that if is compactly assembled and the functor is exact.
We denote by the category of coherently assembled abelian categories, where the -morphisms are strongly continuous exact functors.
Remark 1.42.
Note that a Grothendieck abelian category is compactly assembled if and only if satisfies (AB6) (the same proof as in [E24, Proposition 1.53]).
We recall the following statement from [E24].
Proposition 1.43.
[E24, Proposition E.8] Let be a Grothendieck abelian category which is compactly generated. Then is coherently assembled if and only if it is locally coherent.
For completeness we give the following non-standard characterization of locally noetherian abelian categories.
Proposition 1.44.
Let be a Grothendieck abelian category. The following are equivalent.
is locally noetherian.
is coherently assembled and for any object the map is a monomorphism in
Proof.
. Since is locally noetherian, it is locally coherent, hence coherently assembled by Proposition 1.43. Any object is a directed union of noetherian subobjects hence the map is a monomorphism.
. The assumption implies that the functor
is exact. Moreover, an object is compact if and only if
Now take some and let where is directed. Since the map is a monomorphism, we may and will assume that each map is a monomorphism in For each the map is zero, hence by the left exactness of Therefore, each is compact and is compactly generated. By Proposition 1.43, is locally coherent. It remains to show that the class of compact objects in is closed under subobjects. This again follows from the left exactness of if is compact and then hence i.e. is compact. ∎
We record the special cases of general results on coherently assembled exact categories.
Corollary 1.45.
Let be a coherently assembled abelian category.
is locally -coherent.
The unseparated derived category is dualizable and the functor is strongly continuous. Moreover, we have a well-defined functor
Proof.
The localizing invariants of coherently assembled abelian categories are obtained by applying Definition 1.35.
Proposition 1.46.
Let be a Grothendieck abelian category and let be a coherently assembled abelian category. Suppose that we have a retraction such that both and are left exact and commute with filtered colimits. Then is coherently assembled.
Proof.
The retraction implies that is compactly assembled. The functor is automatically right exact (it is a left adjoint). Hence, we only need to check the left exactness. This follows from the retraction since both and are left exact. ∎
Some non-trivial examples of (not locally coherent) coherently assembled abelian categories are given in Section 8.
We recall the following notation. We denote by the category of Grothendieck abelian categories, where the -morphisms are colimit-preserving left exact functors. We denote by the non-full subcategory with the same objects, where -morphisms are functors such that the right adjoint commutes with filtered colimits. The following statement is certainly known to experts, but we could not find a reference.
Proposition 1.47.
The categories and have filtered colimits, and the (non-full subcategory inclusion) functors commute with filtered colimits.
Proof.
Consider an ind-system By [Lur18, Proposition C.5.5.20] we have a fully faithful functor left adjoint to the functor Putting we obtain an ind-system in such that (functorially in ). By [Lur18, Proposition C.3.3.5] the category has filtered colimits, which are preserved by the functor Recall that a colimit in can be identified with the limit of the associated diagram in (with right adjoint transition functors). The right adjoints preserve discrete objects, which implies that is a Grothendieck abelian category (the former colimit is also computed in ). Using again the fully faithfulness of the functor we conclude that the functor creates filtered colimits, as stated.
Note that is a (strictly) full subcategory. The following result is an (almost) straightforward generalization of the following fact: if is an ind-system of (ordinary) associative coherent rings with left flat transition maps, then the ring is right coherent.
Proposition 1.48.
The full subcategory is closed under filtered colimits.
Proof.
Let be a directed poset, and let be a functor.
First suppose that each is locally coherent. Then we have an -indexed diagram of small abelian categories, The transition functors are exact, hence the colimit is abelian and we have
The general case follows since the diagram in is a retract of the diagram ∎
Finally, we point out that the weak (AB5) axiom holds.
Proposition 1.49.
The categories and satisfy the weak (AB5) axiom: the class of fully faithful functors is closed under filtered colimits.
Proof.
Finally, we mention an expected statement about localizations (Serre quotients).
Proposition 1.50.
Let be a coherently assembled abelian category, and let be a coherently assembled localizing subcategory such that the inclusion functor is strongly continuous. Then the category is coherently assembled and the quotient functor is strongly continuous.
If moreover and are locally coherent, then we have an equivalence
Proof.
The equivalence in the locally coherent case follows from the universal properties.
Denoting by and the right adjoints, for we have a functorial short exact sequence
By assumption, is continuous, hence so is The retraction implies that is coherently assembled by Proposition 1.46. The strong continuity of is already established. ∎
1.9. The situation of dévissage
The conditions of Quillen’s dévissage theorem [Qui73, Theorem 4] naturally generalize to coherently assembled abelian categories. However, this is not immediately straightforward and requires some clarification.
We first treat the case of small categories.
Proposition 1.51.
Let be a small abelian category, and let be a strictly full abelian subcategory such that the inclusion functor is exact. Suppose that every object of has a finite filtration with subquotients in Then is closed under taking subobjects and quotients in
Proof.
It suffices to prove that for any morphism in if then Let be a finite increasing filtration such that and If then there is nothing to prove. If then put and Then and applying induction on we obtain ∎
Next we consider the locally coherent case.
Proposition 1.52.
Let be a small abelian category and a strictly full abelian subcategory such that the inclusion functor is exact. We identify with its essential image in The following are equivalent.
Every object of has a finite filtration with subquotients in
is closed under subobjects and quotients, and every object of has an -indexed increasing exhaustive filtration with subquotients in
generates via extensions and filtered colimits.
Proof.
. By Proposition 1.51, is closed under subobjects and quotients, hence the same holds for Abusing the notation we write for Then the right adjoint commutes with filtered colimits and for any the map is a monomorphism. More precisely, is the largest subobject of which is contained in Define the functorial non-negative filtration inductively: and is the preimage of By induction each of the endofunctors commutes with filtered colimits. It follows from (i) that for any we have for large It follows that for any the filtration is exhaustive. By construction, its subquotients are in which proves the implication.
The implication is evident.
. We identify with the heart of the standard -structure on Then (iii) implies that generates as a localizing subcategory. Hence, generates as an idempotent-complete stable subcategory. Let be the smallest weak Serre subcategory containing i.e. is generated by via extensions. Denote by the (idempotent-complete stable) subcategory consisting of complexes with homology in Then is contained in hence the latter category coincides with and This proves the implication. ∎
Finally, we consider the coherently assembled case.
Proposition 1.53.
Let be a coherently assembled abelian category, and let be a strictly full coherently assembled abelian subcategory such that the inclusion functor is strongly continuous and exact. The following are equivalent.
is closed under subobjects and quotients, and any object of has an -indexed exhaustive filtration with subquotients in
generates via extensions and filtered colimits.
Proof.
Again, the implication is evident.
. Denote by the weak Serre subcategory generated by (note that is not necessarily closed under countable coproducts in ). By Proposition 1.11 the full subcategory is closed under extensions. By assumption the functor sends to Since is exact, the condition (ii) implies that we have
We show that is closed under subobjects, and the assertion about quotients follows. Let and a priori Then hence by Proposition 1.52 (applied to ) we have Applying the colimit functor, we obtain as stated.
The same argument shows the existence of a filtration with required properties. Namely, let then by Proposition 1.52 the object has an -indexed filtration with subquotients in Applying the colimit functor we obtain the required filtration on This proves the implication. ∎
The following is a basic example when the dévissage condition holds.
Proposition 1.54.
Let be a coherenly assembled category. Let be the category of pairs where and is a locally nilpotent endomorphism, i.e. Then is coherently assembled, and the inclusion satisfies the conditions of Proposition 1.53.
Proof.
The only statement requiring a proof is that is coherently assembled. This can be deduced from Proposition 2.25 below. For competeness we give a more direct proof.
For any consider the full subcategory consisting of pairs such that Then we have a direct sequence in By Proposition 1.49, the functor is fully faithful. It is an equivalence since is generated by via filtered colimits. By Proposition 1.48 it suffices to prove that is coherently assembled for each
We first note that is compactly assembled: the forgetful functor commutes with colimits, and the essential image of its left adjoint generates the target via colimits. Next, the right adjoint is continuous, and is exact and conservative, hence is also exact and conservative. This formally implies that is coherently assembled since is. ∎
We mention an example showing that one has to be cautious when dealing with the dévissage condition for coherently assembled categories. Suppose that we have a Grothendieck abelian category and a strictly full Grothendieck abelian subcategory such that the inclusion functor is exact and colimit-preserving, the right adjoint is continuous, is closed under subobjects and quotients in and every object of has an -indexed exhaustive filtration with subquotients in It is tempting to think that if is coherently assembled, then so is However, the following example shows that this is not the case in general, even if is locally coherent.
Example 1.55.
Let be a field, and consider the ring Let be the category of -modules such that acts locally nilpotently on Then is a Grothendieck category. Let Then is locally coherent since is a sequential colimit of noetherian rings with flat transition maps. Further, is closed under subobjects and quotients, and any object has a filtration given by with and The right adjoint to the inclusion is given by hence it is continuous. However, is not coherently assembled.
Assume the contrary. Then by Corollary 1.45 for any compact object the functor commutes with filtered colimits. Clearly, is compact in Consider the direct sequence given by Put Take the element corresponding to the short exact sequence
Then does not factor through for any since is a quotient of This gives a contradiction, hence is not coherently assembled.
2. Compactly assembled -structures on dualizable categories
2.1. Reminder on -structures
We start by recalling the basic notions related to -structures, also fixing the notation. The basic reference is [BBD82] in the context of triangulated categories, and [Lur17a, Section 1.2], [Lur18, Appendix C] in the context of stable categories.
Recall from [BBD82] that a -structure on a triangulated category is a pair of full subcategories such that we have for and for any there exists an exact triangle
A -structure on a stable category is a pair of full subcategories such that is a -structure on We will only deal with -structures on stable categories.
If is another stable category with a -structure then an exact functor is called left -exact if and right -exact if Further, is called -exact if it is both left and right -exact.
Given a -structure on a stable category for any we put
We denote by the right adjoint to the inclusion, and by the left adjoint to the inclusion. We also denote by the same symbols the compositions This should not lead to confusion.
For integers we put In particular, is the heart of the -structure, which is an abelian category. We denote by the functor Since we are working with not necessarily -linear categories, we use the notation
We recall the notation
The -structure is called left bounded resp. right bounded resp. bounded if resp. resp.
Recall that for a -structure on the left completion is given by the limit This is a stable category, and it naturally has a -structure where We have a natural exact and -exact functor If it is an equivalence, then the -structure on is called left complete. Dually, we have a notion of a right completion and similarly for right completeness. Note that in general the right completion of is equivalent to the left completion of and both are given by
It is convenient to use the following terminology for brevity.
Definition 2.1.
A small -category is a small stable category with a bounded -structure
We recall that if is a small abelian category (considered also as an exact category), then we have In particular, if is a small -category, then we have a realization functor corresponding to the inclusion
2.2. -structures compatible with filtered colimits
In this subsection we consider -structures on presentable stable categories, before specializing to dualizable categories.
Recall that an accessible -structure on is said to be compatible with filtered colimits if the subcategory is closed under filtered colimits (note that is automatically closed under all colimits). Equivalently, this means that the functor commutes with filtered colimits. In this case is a Grothendieck prestable category [Lur18, Proposition C.1.4.1]. In particular, the heart is a Grothendieck abelian category.
It is convenient to introduce the following notation.
Definition 2.2.
Let be a presentable stable category, and let be an accessible -structure compatible with filtered colimits. We denote by the localizing subcategory generated by and similarly for
We say that the -structure is continuously bounded if
Note that equivalently is generated by is generated by and is generated by Since the -structure is accessible, each of the categories and is presentable.
We first make some almost tautological observations.
Proposition 2.3.
Let and be as in Definition 2.2. Then the -structure on induces the -structures on and with the same heart. Moreover, we have equalities (of strictly full subcategories of ).
Proof.
The first statement is evident: the (continuous) composition preserves each of the subcategories and Next, we have an inclusion hence On the other hand, is generated via extensions and filtered colimits by hence we have The same argument proves the equality ∎
We recall the following relation with the right completion.
Proposition 2.4.
Let and be as in Definition 2.2.
We have
The right adjoint to the inclusion is given by In particular, it commutes with colimits.
We have if and only if the -structure is right complete.
Proof.
All the assertions are well-known. Namely, the right completion of is identified with see for example [AN21, Lemma A.8]. For convenience, we write the objects of as where and we assume the isomorphisms The natural functor is given by Its left adjoint is given by
Since the -structure is compatible with filtered colimits, the adjunction counit is an isomorphism, i.e. is fully faithful. It follows that the essential image of is identified with This proves all the assertions. ∎
We deduce the following statement, which will have an important -theoretic application.
Proposition 2.5.
Let and be as in Definition 2.2.
The following (lax commutative) square commutes:
(2.1) Here the horizontal functors are the inclusions and the vertical functors are right adjoints to the inclusions.
We have a semi-orthogonal decomposition in
(2.2) In particular, we have an equivalence
Proof.
Next, 2.3 follows directly from 2.1. Namely, the inclusion is colimit-preserving, hence it has a right adjoint. The right orthogonal to its image is identified with or equivalently with It follows from 2.1 that the image of in is contained in the right orthogonal to Since and generate we obtain the semi-orthogonal decomposition (2.2). The equivalence (2.3) follows. ∎
2.3. Compactly assembled -structures
We first recall the notion of a compactly generated -structure.
Proposition 2.6.
Let be a small stable category with a -structure Then the category has a -structure where is the essential image of the functor and similarly for
Moreover, this -structure on is continuously bounded if and only if the -structure is bounded.
Proof.
The first assertion is [Lur18, Lemma C.2.4.3]. For the second assertion, note that the -structure is bounded iff is generated by as an idempotent-complete stable subcategory, which equivalently means that is generated by as a localizing subcategory, i.e. the -structure is continuously bounded. ∎
The following is immediate.
Corollary 2.7.
In the situation of Proposition 2.6, for any integers the full subcategory is the essential image of the functor
Proof.
Clearly, the essential image of is contained in On the other hand, for an object the maps and are isomorphisms. This proves the inverse inclusion. ∎
It is technically convenient to consider the natural -structure on the “twice large” category when is a presentable stable category with an accessible -structure. Namely, is a union of for sufficiently large regular the -structure on induces a -structure on and the inclusion functors are -exact. The compatibility of with filtered colimits exactly means that the functor is -exact (it is automatically right -exact).
We introduce the following natural generalization of a compactly generated -structure.
Definition 2.8.
Let be a dualizable category, and let be an accessible -structure compatible with filtered colimits. We say that this -structure is compactly assembled if the functor is -exact.
We will say that is a dualizable -category if the -structure is compactly assembled and continuously bounded.
Remark 2.9.
In the situation of the Definition 2.8 the functor is automatically right -exact, since its right adjoint is -exact.
A compactly generated -structure is compactly assembled.
Let be a compactly assembled -structure on a dualizable category It follows from Proposition 2.5 that the condition is equivalent to the equalities By Proposition 2.4 the condition is equivalent to the right completeness. One can show that if and only if the Grothendieck prestable category is anticomplete in the sense of [Lur18, Definition C.5.5.4]
To avoid confusion we spell out the following special case.
Proposition 2.10.
Let be a compactly generated presentable stable category, and let be an accessible -structure compatible with filtered colimits. Then this -structure is compactly assembled if and only if it induces a -structure on which equivalently means that we are in the situation of Proposition 2.6.
Proof.
We need to show that for we have This follows from the isomorphisms in
As usual with dualizable categories, we obtain the -accessibility statement.
Proposition 2.11.
Let be a compactly assembled -structure on a dualizable category Then this -structure is -accessible.
Proof.
Let then We obtain hence This proves that the -structure on induces a -structure on Recall that is -presentable by [E24, Corollary 1.21]. Since the -structure on commutes with filtered colimits, it follows that this -structure is -accessible. ∎
Next, we have an interpretation via retracts.
Proposition 2.12.
Let be a dualizable category with a -structure compatible with filtered colimits. The following are equivalent.
The -structure is compactly assembled.
There exists a small stable category with a -structure and a retraction such that both and are colimit-preserving and -exact.
There exists a dualizable category with a compactly assembled -structure and a retraction such that both and are colimit-preserving and left -exact.
Moreover, similar assertions are equivalent for compactly assembled continuously bounded -structures, where in (ii) the -structure is required to be bounded.
Proof.
. By Proposition 2.11 it suffices to take If the -structure is continuously bounded, then we instead take Note that the functor indeed sends to since it sends to
is trivial.
. By Remark 2.9 the functor is automatically right -exact, so we only need to show the left -exactness. This follows from the retraction, since is identified with the composition of left -exact functors
Suppose that the -structure on is continuously bounded. We need to show that the -structure on is also continuously bounded. By Remark 2.9 it suffices to show the equalities First, is generated by hence It remains to show that Take and put Then hence We get so This proves the proposition.
Corollary 2.13.
Let be a dualizable category with a compactly assembled -structure Let be a dualizable subcategory such that the inclusion functor is strongly continuous. Suppose that the -structure on induces a -structure on Then the latter -structure is also compactly assembled.
If moreover the -structure is continuously bounded, then so is the -structure
Proof.
Denote by the (continuous) right adjoint to Then is left -exact and is -exact. Hence, the retraction implies that is compactly assembled by Proposition 2.12. The “moreover” assertion also follows from loc. cit. ∎
It is useful to keep in mind a reformulation in terms of compact morphisms (although we will not use it).
Proposition 2.14.
Let be a dualizable category, and let be an accessible -structure compatible with filtered colimits. The following are equivalent.
The -structure is compactly assembled.
For any compact morphism in the morphism is also compact in
Proof.
This is formal: the -structure is compactly assembled if and only if the endofunctor is strongly continuous. For completeness we give the details.
. Compactness of means that factors through in Since we conclude that is also a compact morphism.
. Consider as a continuous endofunctor of and similarly for We need to show that the morphism is an isomorphism of functors from to Let and where is directed. Then for each there exists such that the transition morphism is compact, hence so is This proves that the map is an isomorphism, as required. ∎
The following basic properties of the categories are almost immediate.
Proposition 2.15.
Let be a dualizable category with a compactly assembled -structure
Let where The exact category (with the induced exact structure from ) is coherently assembled and the inclusion functor is strongly continuous. In particular the abelian category is coherently assembled.
The categories and are dualizable, and their inclusion functors into are strongly continuous. The induced -structures are compactly assembled.
Proof.
From now on, in the situation of Proposition 2.15 (i) we write for Also, for a morphism in the compactness of in is equivalent to its compactness in which we will tacitly use in the proofs. Note that for the inclusion functor is exact and strongly continuous.
We formulate the following statement only for dualizable -categories, but it holds in a much more general setting, see Remark 2.17 below.
Proposition 2.16.
Let be a dualizable -category (Definition 2.8). Then we have an equivalence
Proof.
By Proposition 2.12 we have a retraction where is equipped with a bounded -structure, and the functors and are continuous, exact and -exact. Then the functor is a retract of the functor (in ), hence we may and will assume that
Remark 2.17.
The statement of Proposition 2.16 holds more generally when is presentable stable, the accessible -structure is right complete and compatible with filtered colimits, and the Grothendieck prestable category is anticomplete.
The following two important statements are special cases of much more general results on Grothendieck abelian -categories, but our proofs use only the general machinery of presentable stable envelopes of coherently assembled exact categories.
Proposition 2.18.
Let be a dualizable -category, and let be an integer. Denote by the universal functor. Then for the map
is an isomorphism for and a monomorphism for
Proof.
We have isomorphisms
Applying Corollary 1.38, we see that the map
is an isomorphism for and a monomorphism for as required. ∎
Recall that in the situation of Proposition 2.16 we have – the unseparated derived category of the heart.
Proposition 2.19.
Let be a dualizable -category and let be an integer. Consider the (strongly continuous exact) functor induced by the inclusion The following are equivalent.
is an equivalence.
For the map is an isomorphism for and a monomorphism for
For the map is an isomorphism for
Proof.
The implication follows directly from Proposition 2.18.
The implication is trivial.
. Clearly, the essential image of generates as a localizing subcategory. Hence, we only need to show that is fully faithful. By Proposition 1.39, it suffices to show that for there exists and a map such that the map is an epimorphism. If then there is nothing to prove. If consider the map It follows from (iii) (applied to ) that we can find an epimorphism in such that the composition is zero. Then factors through so we obtain a map Continuing the process, we find an epimorphism in such that the composition is zero. This means that factors through which proves the implication. ∎
We conclude this subsection with a -theoretic application of the above results.
Corollary 2.20.
Let be a dualizable category with a compactly assembled -structure Then the following square of spectra is cartesian:
2.4. General constructions of -structures
We will need a general efficient method for constructing a -structure. First we deal with small categories, for which the construction seems to be a folklore knowledge. In the following proposition we do not a priori assume the idempotent-completeness, but it holds a posteriori.
Proposition 2.21.
Let be a small stable category, and let be a small abelian category. Suppose that is a fully faithful exact functor such that generates as a stable subcategory. Then has a unique bounded -structure such that Each object of has a finite filtration with subquotients in
Moreover, if induces isomorphisms for then we have
Proof.
We will construct the -structure with required properties, and its uniqueness follows from the proof.
We define resp. to be the full subcategory generated via extensions by where resp. Clearly, and Also, the orthogonality holds since for by assumption.
It remains to prove that for any there exists some and a functor such that and we have for where If this is the case for some object we say that has a good filtration. Since generates as a stable subcategory, we only need to show that for any object with a good filtration, for any for any and for any morphism the object has a good filtration.
Since the class of objects with a good filtration is closed under shifts, we may and will assume that Let and be as above, with etc. If then and there is nothing to prove. So we assume that If then and has a good filtration.. Hence, we assume that We take the largest such that then we get a unique factorization It suffices to prove that has a good filtration with subquotients in We prove this by induction on
If then hence the cofiber sequence shows that has a good filtration with subquotients in Now suppose that and the assertion holds for smaller values of We have so there is some and an isomorphism Denote by the morphism in such that the composition is homotopic to First suppose that is not an epimorphism. Consider the composition and put Then we have a unique factorization and it suffices to prove that has a good filtration with subquotients in Hence, we may and will assume that is an epimorphism. Put then the composition factors uniquely through Moreover, we have By the induction hypothesis, this object has a good filtration with subquotients in This proves the induction step.
Finally, the “moreover” assertion is immediate: the assumption implies that is closed under extensions in hence the functor is essentially surjective. ∎
We deduce an analogue for dualizable categories.
Proposition 2.22.
Let be a dualizable category, and let be a coherently assembled abelian category. Suppose that is a strongly continuous fully faithful exact functor, such that generates as a localizing subcategory.
has a unique compactly assembled continuously bounded -structure such that Moreover, generates via extensions and filtered colimits.
If induces isomorphisms for then we have
Let be another dualizable -category and let be a strongly continuous exact functor. Then is -exact if and only if we have
Proof.
(i) Again, we construct the -structure with required properties, and its uniqueness follows automatically.
Denote by the (small) stable subcategory generated by By Proposition 2.21, has a unique bounded -structure such that The strong continuity of implies that Since generates as a localizing subcategory, it follows that
We claim that the -structure on induces a -structure on To see this, it suffices to prove that the composition is -exact. Since the -structure on is bounded, we only need to show the inclusion By Proposition 2.21, each object of has a finite filtration with subquotients in Hence, it suffices to show that The latter inclusion follows from the strong continuity of
We claim that the constructed -structure on satisfies the desired properties. Since the -structure on is compactly assembled and continuously bounded, so is the -structure on by Corollary 2.13. Since generates via extensions (by Proposition 2.21), the retraction implies that is generated by via extensions and filtered colimits.
(ii) follows from the proof of (i): we have by Proposition 2.21. Hence, is generated by via filtered colimits, which shows that the functor is essentially surjective.
(iii) The “only if” direction is clear. For the “if” direction we first note that since is generated by via extensions and filtered colimits. Next, we have since generates via extensions and colimits. Finally, we have since is generated by via finite limits, extensions and filtered colimits (by the right completeness). ∎
Corollary 2.23.
Let be dualizable -category and let be an integer. There is a unique compactly assembled continuously bounded -structure on such that the functor is -exact. Moreover, we have
Proof.
This follows directly from Proposition 2.22, applied to the functor ∎
Let be a dualizable category and let be a continuous exact monad on i.e. Here we consider the category as a presentable stable -monoidal category (i.e. an -algebra in ). Then we can consider the dualizable category of -modules in The forgetful functor has a left adjoint which is automatically strongly continuous. By [E24, Proposition C.1] this gives a fully faithful functor
Its essential image consists of pairs such that generates as a localizing subcategory. For such the functor is monadic, so we have an equivalence
More generally, The right adjoint to (2.6) is given by
We now apply this point of view to dualizable -categories.
Corollary 2.24.
Let be a dualizable -category. Let be a continuous exact monad. Suppose that the functor is left -exact. Then the category has a unique compactly assembled continuously bounded -structure such that the functor is -exact.
Proof.
By adjunction, the left -exactness of the functor implies that the composition is fully faithful. Since its image generates the target as a localizing subcategory, the assertion is a special case of Proposition 2.22. ∎
We also deduce the following basic result.
Proposition 2.25.
Let be a dualizable -category. Let be a formal variable of degree and consider the (compactly generated) category of -modules such that acts locally nilpotently on the homotopy groups. Denote by the restriction of scalars functor for
The category has a unique compactly assembled continuously bounded -structure such that the functor is -exact. Moreover, the abelian category is identified with the category of pairs where and is a locally nilpotent endomorphism.
If we have an equivalence then we also have an equivalence
Proof.
(i) It is well-known that the category is generated by the single compact object in particular the functor is strongly continuous. Thus, the functor is also strongly continuous and its image generates the target as a localizing subcategory.
We have in particular the functor is left -exact. Applying Corollary 2.24, we obtain the stated -structure and its uniqueness.
By construction the forgetful functor is -exact and conservative. This gives the description of the heart and proves (i).
(ii) By Proposition 1.31 and its proof, we may and will assume that is locally coherent, i.e. is compactly generated. Put and let be the category of pairs where and is a nilpotent endomorphism. We need to prove that the realization functor
is an equivalence. By [Nee21, Lemma 2.1] it suffices to show that for there exists and a map such that Denote by the forgetful functor. Then there exists some and a map such that By adjunction corresponds to a map where we consider as an object of Choose such that is acting by zero on Then factors through (non-uniquely), which gives a map such that is an epimorphism, i.e. Since this shows that (2.7) is an equivalence and proves (ii). ∎
2.5. Sufficient conditions for left -exactness
In this subsection we consider dualizable -categories, and we record some semi-trivial observations on how to check the left -exactness of a functor between them. The statements of course hold in more general situations.
Proposition 2.26.
Let be a continuous exact functor between dualizable -categories. The following are equivalent.
is left -exact.
We have
Proof.
The -structure on is right complete and -accessible, hence is generated by via finite limits, extensions and filtered colimits. This implies the equivalence. ∎
Proposition 2.27.
Let and be dualizable -categories, and let be a strongly continuous, exact and -exact functor, such that generates as a localizing subcategory and moreover the restriction is fully faithful. For a continuous exact endofunctor the following are equivalent.
is left -exact.
The composition is left -exact.
Proof.
The implication is trivial since both and are left -exact.
. We first show that is left -exact. By Proposition 2.22, generates via extensions and filtered colimits. Hence, Thus, by Proposition 2.26 the functor is left -exact.
Now we show that is left -exact. Let By adjunction for we have Again, since generates via extensions and filtered colimits, for any we have But generates via colimits and extensions, since We conclude that as required. ∎
2.6. Filtered colimits of dualizable -categories
The following result is elementary, and again it is a special case of a more general fact.
Proposition 2.28.
Let be a directed system of dualizable -categories, where we assume the transition functors to be strongly continuous, exact and -exact. Then the colimit is naturally a dualizable -category and all the functors are -exact.
Proof.
If all are compactly generated, then the assertion follows directly from Proposition 2.10: we have a diagram of small -categories and the colimit (taken in ) has a natural bounded -structure such that all the functors are -exact. We have which proves the proposition in this case.
The general case follows: take the colimit in and let be the colimit (in ) of functors Then is fully faithful and the composition is -exact, hence by Corollary 2.13 we obtain a compactly assembled continuously bounded -structure on The functors are -exact since they are compositions of -exact functors ∎
We will need the following elementary application.
Proposition 2.29.
Let be a dualizable -category. Let be a poset, and consider the -structure on the (dualizable) category given by
Then this -structure is accessible, compatible with filtered colimits and is continuously bounded. Moreover, this -structure is compactly assembled in the following two cases:
is finite.
is a meet-semi-lattice, i.e. has the smallest element and for any there exists (in other words, considered as a category has finite products).
Proof.
It is clear that the -structure is well-defined, accessible and compatible with filtered colimits. Denote by the left adjoint to Then the functors form a conservative family, hence is generated by so the -structure is continuously bounded. Now consider the cases.
(i) By Proposition 2.12 we may and will assume that is compactly generated. Then is also compactly generated and we have since is finite. It follows that the -structure on induces a -structure on which proves (i).
(ii) Let be the poset of (full) finite subposets closed under finite meets (in particular, we require that contains the smallest element of ). Then is directed, and we have where the colimit is taken in (equivalently, in ). Here the transition functors are left Kan extensions. Now, each inclusion of posets in has a left adjoint by assumption. Hence, the functor is -exact. It remains to apply (i) and Proposition 2.28. ∎
2.7. Dual -structures
This subsection can be skipped on the first reading, we include it for completeness. Recall that if is a small stable category with a -structure then the category has a natural opposite -structure given by By taking ind-objects we obtain the notion of a dual of a compactly generated -structure. We explain a natural generalization for compactly assembled -structures.
Proposition 2.30.
Let be a dualizable category with a compactly assembled -structure Then the category has a natural compactly assembled -structure which is described as follows. Consider the (strongly continuous fully faithful) functor Then
Proof.
Consider the compactly generated -structure on coming from the opposite -structure on It suffices to prove that it induces the -structure on via which would be automatically compactly assembled by Corollary 2.13. Indeed, by definition is the preimage of and the condition means that the mapping space is discrete for or equivalently
Let It suffices to prove that the object is contained in the essential image of If we identify with then Consider the functor
Then commutes with finite limits (since finite limits in are the same as in ). Hence, is pro-representable, i.e. there exists a directed poset and a functor such that Then
The essential image of in is identified with the full subcategory Hence, we need to prove that the functor commutes with countable colimits. It suffices to show that for any there exists such that the map is compact in hence also compact in This follows directly from the fact that the functor (2.8) commutes with filtered colimits. Namely, it suffices to show that for and for there exists a compact map and an element such that It suffices to take for large where This proves that as required. ∎
The above construction is functorial in the following sense.
Proposition 2.31.
Let and be dualizable categories with compactly assembled continuously bounded -structures resp. Let be a strongly continuous, exact and -exact functor. We equip and with -structures as in Proposition 2.30. Then the functor is -exact.
Proof.
We have a commutative square
in which the functors are strongly continuous and the horizontal arrows are fully faithful. Clearly, the right vertical arrow is -exact. By the proof of Proposition 2.30 the horizontal arrows are -exact. Hence, the left vertical arrow is also -exact. ∎
Remark 2.32.
It is easy to show that in the situation of Proposition 2.30 the double dual -structure on is identified with the original -structure Furthermore, the -structure is continuously bounded if and only if so is This is also equivalent to the right completeness of both -structures.
3. Refined -theoretic theorem of the heart for dualizable categories
In this section we prove the following result.
Theorem 3.1.
Let be a dualizable category with a compactly assembled continuously bounded -structure Let be an integer. Then the map is an isomorphism for and a monomorphism for
We first mention an immediate corollary.
Corollary 3.2.
Let be a dualizable -category. Suppose that for some we have isomorphisms for and (this assumption always holds for ) Then the map is an isomorphism for and a monomorphism for
Proof.
Our arguments are completely different from [Bar15]. Even if we are only interested in still the the proof of [Bar15, Theorem 6.1] does not generalize to dualizable (not compactly generated) categories. We also note that an example below (Corollary 8.2) shows that the group might be non-zero for a dualizable -category. In other words, the statement of [AGH19, Theorem 2.35] does not hold for dualizable categories. However, a different version of a construction from loc. cit. will be very useful for us, see Propositions 3.4 and 3.5 below.
One of the key statements for our proof is the following special case of Theorem 3.1.
Lemma 3.3.
Let be a dualizable -category. Then for any the map is surjective.
The proof of this lemma will require some preparations. The general idea is to find “sufficiently bounded” representatives of classes in the negative continuous -theory of The key ingredient is the following construction. We consider the set as a totally ordered set with the usual order.
Proposition 3.4.
Let be as in Theorem 3.1. Consider the small stable category with the bounded -structure induced from and let be its simultaneous left and right completion, i.e. . Equivalently, is the left completion of the category i.e.
Consider the localizing subcategory generated by objects of the form where is a functor such that for any rational and for any the morphism is compact in Here we identify with
The category is dualizable and the inclusion functor is strongly continuous. Moreover, the above generating collection of objects of is exactly the collection of all -compact objects of up to isomorphism.
The (compactly generated) -structure on induces a -structure on The latter -structure is compactly assembled. The category consists of objects isomorphic to where is a functor such that for any rational the map is compact in (equivalently, compact in ).
We have a -exact equivalence so that the following square commutes:
Here the lower horizontal functor is induced by the inclusion
Proof.
(i) The argument is the same as for similar constructions with dualizable categories, but we spell it out for completeness. Denote by the full subcategory formed by objects isomorphic to an object in the generating collection described above. Take some where is a functor as above. Then for any the object is also in and we have Moreover, the map in factors through the constant ind-object hence this map is compact in This implies that is dualizable and the inclusion functor is strongly continuous, see also [E24, Lemma1.83 and its proof].
To prove the equality (of strictly full subcategories of ), we only need to show that is closed under countable colimits and shifts. Clearly, for we have for Next, let be a collection of functors as above. For define to be the left Kan extension of i.e.
Then for any there are only finitely many such that hence the direct sum is well-defined. Moreover, for any rational and for the map is compact in We have
which shows that is closed under countable coproducts.
It remains to show that is closed under cones (cofibers). Let be a morphism in and where the functors satisfy the above conditions. We may assume that is induced by a morphism of functors Indeed, identifying resp. with resp. we first see that there exists a strictly increasing function such that is induced by a morphism of functors Take any strictly increasing function such that for We obtain a morphism of functors (which factors through the right Kan extension of ), which induces Replacing by we may assume that we have a morphism which induces
Now we have It suffices to check that for and for the map is compact in By assumptions on and the maps
are compact in It follows that the map
is also compact in Applying to (3.1) we deduce that the map is compact in This finishes the proof of (i).
Now (ii) follows directly from (i) and Corollary 2.13. Indeed, it suffices to show that for we have where is the truncation endofunctor of By (i) we have where satisfies the above conditions. Then the functor also satisfies the above conditions, hence
We also have
which gives the stated description of This proves (ii).
Finally, we prove (iii). Consider the composition Clearly, is strongly continuous and fully faithful. It suffices to show that the essential image of coincides with Take some object and let where and each map is compact in Arguing as in [E24, Proof of Proposition 1.82] we can extend the functor to a functor such that for any rational the map is compact in Then we have where we identify with In particular, we have and by (ii) any isomorphism class in can be obtained in this way. This proves (iii). ∎
The following is a slightly more sophisticated version of Eilenberg swindle.
Proposition 3.5.
Let be as in Proposition 3.4. Then any localizing invariant vanishes on and In particular, we have We obtain an isomorphism
Proof.
It suffices to show that we have the vanishings
Note that is contained in as a subcategory of and the inclusion functor is strongly continuous. For an object the coproduct exists in Hence, we have a strongly continuous endofunctor
We claim that preserves the subcategory It suffices to prove that for we have By Proposition 3.4 we have where is a functor such that for any rational and for any the morphism is compact in Define by the formula
Then for and we have
Hence, for rational the map is compact. Therefore, as required.
Denote by the restriction of Then is strongly continuous and by construction we have This proves the first vanishing in (3.3). The other vanishing is proved by a similar argument.∎
We will need to consider the natural -structures on the iterated Calkin categories. Recall from Subsection 1.6 that for a dualizable category we denote by the compactly generated category We will denote by the -th iteration of this construction, It is convenient to deal with them inductively, using the following observation.
Proposition 3.6.
Let be a dualizable category with an -accessible -structure compatible with filtered colimits. Suppose that for some the functor is left -exact. Consider the functor
The -structure on induces an -accessible -structure on via the inclusion
Moreover, this -structure is compatible with filtered colimits.
The functor is right -exact, and the functor is left -exact.
For the functor is left -exact.
Proof.
(i) is straightforward since the colimit functor is -exact. Note that the -accessibility of the induced -structure on follows from the identification of with the kernel of the -exact functor
To prove (ii), it suffices to consider the restriction of to Considering as a full subcategory of we have
since the functor is left -exact by assumption. This proves that the functor is left -exact, as stated. It is also clear that is right -exact.
We prove (iii). Consider again as a full subcategory of and recall that for the object is compact. Now take any object where and Consider the category as a full subcategory of Then we have
because again by assumption we have ∎
In the following proof we use the notation for where is a small category. We write for the -th iteration of this construction, and for convenience we put
Proof of Lemma 3.3.
Let and be as in Proposition 3.4. By loc. cit. we can identify with and with By Proposition 3.5 we have the following isomorphisms:
It suffices to prove that the image of (the isomorphism) is contained in the image of the map
We identify with Below for any we identify the category with its essential image in and similarly for and More precisely, is the intersection of kernels of the functors
Take some class and choose an object such that Consider the natural -structure on which is obtained by an iterated application of Proposition 3.6. Then the object is a lift of the image of in The latter object represents the element
We can now describe a representative for Namely, consider the composition
Then the object is contained in the full subcategory We then have
Now, by (an iterated application of) Proposition 3.6 the functor is right -exact and the functor is left -exact. In particular, we have On the other hand, since we obtain
Therefore, we have
It follows formally that the image of in is contained in Denote by the image of under the functor
By construction, we have
We claim that is compact. Equivalently, this means that is annihilated by the functor
To see this, consider the similar functor
By Proposition 3.6, both and preserve bounded objects. We obtain a commutative square
Note that is -exact and it induces an equivalence between the hearts. Indeed, the hearts of both the source and the target are identified with the full subcategory of which is the intersection of the kernels of the functors
It follows that is conservative. Now we have and hence Therefore, i.e. as stated.
We conclude that the class is the image of Since was arbitrary, this proves the lemma. ∎
Proposition 3.7.
Let be as in Theorem 3.1 and let be an integer. Consider the exact category whose underlying category is and a morphism is an inclusion in if and only if both morphisms and are monomorphisms in
is a coherently assembled exact category. We denote by the universal continuous exact functor.
We have strongly continuous fully faithful exact functors given by They induce fully faithful functors on presentable stable envelopes, which we denote by the same symbols. They give a semi-orthogonal decomposition in
(3.5) -
We have a short exact sequence in
Here is induced by the identity functor and is the kernel of The functor
is strongly continuous and fully faithful, and its essential image is the heart of a compactly assembled continuously bounded -structure. Moreover, the maps are isomorphisms for where
Proof.
The proof of (i) is straightforward: it is clear that the exact structure is well-defined, it induces an exact structure on and the functors and are exact.
We prove 3.5. Both functors and are exact. The left adjoint to is given by hence it is exact. The right adjoint to is given by hence it is also exact. We have the isomorphisms and the vanishing Moreover, we have a functorial short exact sequence
Passing to presentable stable envelopes we deduce the fully faithfulness of and and obtain the semi-orthogonal decomposition (3.5). This proves 3.5.
Finally we prove (iii). It is clear that is a strongly continuous quotient functor. We first show that (i.e. the kernel of ) is generated by the essential image of as a localizing subcategory. Denote by the additive category with a split exact structure. Clearly, is generated by the images of objects of the form where is an exact inclusion in which means that is a monomorphism in (the notation is introduced in (1.1)). Denote by the image of in It suffices to show that we have an isomorphism where First we observe that is also an exact inclusion in and Hence, we may and will assume that Put then factors uniquely through and the induced morphism is an exact inclusion in Hence as stated.
It is clear that is a strongly continuous exact functor (by definition of the exact category ). Hence, by Proposition 1.34 induces a strongly continuous exact functor By Proposition 2.22 it remains to prove that the induced maps are isomorphisms for and (where can be strictly negative). We take such and denote by the universal functor. By 3.5 we have a cartesian square of spectra:
We first deal with the relevant homotopy groups of the bottom right spectrum. The fully faithfulness of the functor gives the isomorphisms
Next, for we compute:
Therefore, by Proposition 2.18 the right vertical arrow in (3.6) induces isomorphisms on for and a monomorphism for Hence, the same holds for the left vertical arrow in (3.6). But it has a right inverse, induced by the functor We conclude that the map is an isomorphism for as required. This proves (iii). ∎
Proof of Theorem 3.1.
We start with the key cofiber sequence of spectra, which will be used repeatedly. Let and consider the exact category and the dualizable category from Proposition 3.7. By loc. cit. we have a cofiber sequence
Next, we prove that for the map is an isomorphism for each and a monomorphism for By Proposition 3.7 the category from (3.7) has a compactly assembled continuously bounded s-structure, and we have an equivalence By the assertion on from loc. cit. and by Proposition 2.19 we have an equivalence for Hence, by Lemma 3.3 we have a surjection for On the other hand, the composition is left inverse to the realization functor. Therefore the map from (3.7) induces an isomorphism for The long exact sequence of homotopy groups for (3.7) shows that the map is an isomorphism for and a monomorphism for as stated.
Passing to the colimit over and using Proposition 2.16 we deduce that for the map is an isomorphism for and a monomorphism for
It remains to prove that we also have isomorphisms for For this we may and will assume that i.e. we will prove that the map is an isomorphism for We will use the already established case as a base of induction on
Suppose that the statement is proven for where We need to prove that the map is an isomorphism. Take some It follows from the induction hypothesis that in (3.7) the map is an isomorphism (it has a right inverse, which is an isomorphism). It follows from the long exact sequence of homotopy groups that the images of the maps and generate the target. But the image of the latter map is contained in the image of the former map, hence the former map is surjective. Passing to the colimit over and using Proposition 2.16 again, we conclude that the map is surjective. Now let again be a positive integer. Applying the above surjectivity to we see that the realization functor induces a surjection Therefore, its left inverse is an isomorphism. Applying again the long exact sequence for (3.7) we deduce that the map is an isomorphism. Passing to the colimit over we conclude that the map is an isomorphism. This proves the induction step and the theorem. ∎
4. Coconnectivity estimates
4.1. Estimates for -groups
Using Theorem 3.1 we prove the following key result, which is a dual analogue of Waldhausen’s connectivity estimates.
Theorem 4.1.
Let and be dualizable -categories, and let be an integer. Suppose that is a functor satisfying the following conditions:
is strongly continuous, exact and -exact;
generates as a localizing subcategory;
for the map is an isomorphism for and a monomorphism for
Then the induced map is an isomorphism for and a monomorphism for
It is convenient to introduce the following notion.
Definition 4.2.
Remark 4.3.
It is easy to see that for a -coconnective functor the condition on maps between groups automatically holds for all On the other hand, if is compactly generated, then it suffices to check this condition for These assertions follow for example from Proposition 4.4 below.
We first interpret the -coconnectivity condition via monads.
Proposition 4.4.
Let be a dualizable -category. Let be a continuous exact monad. Suppose that the functor is left -exact, so by Proposition 2.24 there is a unique compactly assembled continuously bounded -structure on such that the natural functor is -exact. For the following are equivalent.
The functor is -coconnective.
The functor is left -exact.
Proof.
The implication is immediate by adjunction.
. Again, by adjunction the functor takes to Here we of course used that is generated by via extensions and colimits. By Proposition 2.26 we conclude that this functor is left -exact. ∎
Corollary 4.5.
Let be a -coconnective functor between dualizable -categories, where Then induces an equivalence of exact categories
Proof.
By Proposition 2.22 (ii) the functor induces an equivalence By Proposition 4.4 the functor takes to hence it takes to It follows by adjunction that the functor is fully faithful, and it induces isomorphisms on To see that it is essentially surjective, note that its essential image is closed under extensions and filtered colimits, and it contains for Therefore, this essential image coincides with ∎
We now recall the machinery of deformed tensor algebras, which will be crucial for our arguments. We refer to [E25c, Section 1.9] for this notion in an abstract setting. Here we need the following special case. Let be a dualizable category. We consider the functor
which is left adjoint to the composition
We will typically write assuming that is a continuous exact functor with a chosen morphism Note that by adjunction we have a natural morphism of functors
We consider as a continuous exact monad on
We will use the natural non-negative multiplicative filtration on as explained in [E25c, Section 1.9]. Its associated graded is the (non-deformed) tensor algebra Note that we have and
To deal with monads which are deformed tensor algebras we will use the key short exact sequence given by the following proposition. We recall the following notation: for an accessible exact functor between presentable stable categories we denote by the category of triples where and By [E24, Proposition 1.2] this category is also presentable stable. Moreover, if and are dualizable and is continuous, then the category is also dualizable by [E24, Proposition 1.78].
Proposition 4.6.
Let be a dualizable category, and let be a continuous exact functor with a morphism
We have a short exact sequence in
Here is generated as a localizing subcategory by the image of the strongly continuous functor
(4.1) We consider as a (strongly continuous) functor from to
We denote by the (strongly continuous) inclusion functors
and denote by resp. the (strongly continuous) left adjoint to resp. right adjoint to Then the functors are isomorphic, and they are left adjoint to the forgetful functor. The functors and are left inverse to the functor
The adjunction counit is a split monomorphism, its left inverse is given by
(4.2) This gives a direct sum decomposition in
Proof.
We prove 4.1 using the universal property of the monad First, by [E24, Proposition 1.78] the functors and give a semi-orthogonal decomposition in i.e. By [E24, Proposition 1.79] the functor from (4.1) is strongly continuous. Hence, by [E24, Corollary 1.57] the category (as defined in 4.1) is dualizable and the inclusion functor is strongly continuous. We put and consider the (strongly continuous) composition
By construction, is dualizable, the functor is strongly continuous and generates as a localizing subcategory (since and generate ). Hence, is contained in the essential image of (2.6). For any other pair using [E24, Proposition 1.79] we obtain a natural equivalence of spaces:
Hence, the fully faithfulness of (2.6) implies an equivalence
This proves 4.1 and also the first assertion of (ii). Namely, we see that is left adjoint to the forgetful functor, and the compositions and are isomorphic by the construction of The second assertion of (ii) follows from the definitions.
We apply this general construction to dualizable -categories.
Proposition 4.7.
Let and be as in Proposition 4.6, and we keep the notation from therein. Suppose that is a compactly assembled continuously bounded -structure on Suppose that for some the functor is left -exact. Then the categories and have natural compactly assembled continuously bounded -structures such that the functor is -coconnective and the functor is -coconnective.
Proof.
This follows almost directly from Propositions 4.4 and 4.6. Namely, we need to see that the corresponding monads on satisfy the required assumptions. This is clear for since we have Next, the endofunctor has an exhaustive increasing filtration with subquotients By assumption, is left -exact, hence is left -exact for We conclude that is left -exact. This proves the proposition. ∎
The following general construction gives a way to reduce certain questions about (continuous exact) monads to the case of a deformed tensor algebra.
Proposition 4.8.
Let be a strongly continuous exact functor between dualizable categories, and suppose that generates as a localizing subcategory. We define a functor inductively as follows. We put and for we define
The functor is induced by the natural map of monads on namely
Then we have an equivalence
Proof.
This is in fact the same construction as in [E25c, Proof of Proposition 3.15] if we spell it out in terms of monads on We give the details for completeness.
Denote the transition functors by By construction, generates for all Next, for the morphism of functors factors through Indeed, by construction the monad is the deformed tensor algebra of The term of the standard filtration is identified with Composing with and precomposing with we get a factorization in the category
Moreover, the composition
is homotopic to the identity. Hence, we get an isomorphism of monads
By assumption, generates hence we obtain an equivalence This proves the proposition. ∎
Again, we apply this abstract result to dualizable -categories.
Proposition 4.9.
We keep the notation and assumption from Proposition 4.8. Suppose that and are moreover dualizable -categories and the functor is -coconnective for some There are compactly assembled continuously bounded -structures on the categories which are uniquely determined by the requirement that all the transition functors are -exact. Moreover, for each the functor is left -exact.
Proof.
We construct the -structures inductively, simultaneously proving the left -exactness assertion. We are already given the -structure on By assumption on and by Proposition 4.4, the functor is left -exact.
Suppose that for some we already have the claimed -structures on for such that for these the functor is left -exact. Applying Proposition 4.7, we obtain a unique compactly assembled continuously bounded -structure on such that the transition functor is -exact. We need to prove that the functor
is left -exact. By Proposition 2.27 this is equivalent to the left -exactness of the functor By construction, the monad has a non-negative increasing filtration such that The inclusion of is a right inverse to the morphism Hence, it suffices to show that for the functor is left -exact. This follows from the induction hypothesis: we have
and the latter functor is left -exact since is left -exact and This proves the induction step and the proposition. ∎
Proof of Theorem 4.1.
We first consider the following special case. Let be as in the theorem, and suppose that moreover the monad is isomorphic to where is equipped with a morphism and the functor is left -exact. Put Applying the construction from Proposition 4.6 (and changing the notation for the categories), we obtain a short exact sequence in
As above, we denote by the natural inclusions, and consider the adjoints and Applying to (4.3), we obtain the cofiber sequence
This gives an equivalence of spectra
Now, recall that by Proposition 4.6 we have a functor right inverse to Moreover, by Proposition 4.7 is naturally a -category and the functor is -coconnective. By Corollary 4.5 induces an equivalence By Theorem 3.1 the latter equivalence of exact categories implies that induces an equivalence Hence the left inverse functor also induces an equivalence on We conclude from (4.4) that
Now consider the general situation of the theorem. By Proposition 4.9 we have a sequence of dualizable -categories such that each transition functor is -exact, is generated by and the monad is the deformed tensor algebra of a functor such that is left -exact. By the above special case, we know that the spectrum is -coconnective for Hence, also the spectra are -coconnective. Taking the colimit over we conclude that is -coconnective. This proves the theorem. ∎
4.2. Estimates for higher nil groups
For the proof of the theorem of the heart for it will be important to have a version of Theorem 4.1 for higher nil groups We will need the following immediate application of Proposition 2.25.
Proposition 4.10.
Let and be a dualizable -categories, and let be a -coconnective functor for some Let be a formal variable of degree Then the functor
is also -coconnective, where the source and the target are equipped with -structures from Proposition 2.25.
Proof.
By Proposition 4.4, the -coconnectivity of means that the functor is left -exact, and we need to prove that the same holds for the adjunction counit of the functor (4.5). As in Proposition 2.25 we consider the functor (restriction of scalars for ). Consider the composition
By Proposition 2.27 it suffices to show that the functor is left -exact. This follows from the isomorphism
We recall that for and for the higher nil spectrum is defined as where we take the reduced part in the simplicial sense and we consider the assignment as a simplicial -ring, see [E25c, Section 8]. On the level of homotopy groups we have
As usual, we denote by the associated localizing invariant of dualizable categories.
Corollary 4.11.
Let be a -coconnective functor between dualizable -categories, where Let be an integer. Then the induced map is an isomorphism for and a monomorphism for
Proof.
Recall that we have the identification of reduced localizing motives which follows from the Beilinson’s full exceptional collection on [Bei78], see for example [E25c, Proof of Theorem 9.1]. This isomorphism implies that is naturally a direct summand of and similarly for Hence, the spectrum is a direct summand of the spectrum
By the inductive application of Propositions 2.25 and 4.10, we see that the tensor products and are naturally -categories, and the functor
is -coconnective. Applying Theorem 4.1 again, we see that the spectrum (4.6) is -coconnective. ∎
We record another immediate application of the above results.
Corollary 4.12.
Let be a dualizable -category. Then we have
Proof.
5. Theorem of the heart for
We prove the following result.
Theorem 5.1.
Let be a dualizable -category. Then the realization functor induces an equivalence of spectra
To prove Theorem 5.1 we will deal with the natural filtration on For we consider the localizing invariant i.e.
These functors form a direct sequence We have
and
For convenience we put
Throughout this section we will systematically use the notion of a -coconnective functor from Definition 4.2. Our first step is to deduce the coconnectivity estimates for where is -coconnective for some
Corollary 5.2.
Let be a -coconnective functor between dualizable -categories, where Then for the spectrum is -coconnective.
Proof.
To prove Theorem 5.1 we will consider certain statements depending on four integer parameters (we do not claim that they hold for all values of the parameters, this is certainly not the case). Namely, for and consider the statements
If is a -connective functor between dualizable -categories, then the map of spectra
is null-homotopic.
Let be a dualizable -category, and suppose that is a continuous exact endofunctor with a map such that is left -exact. Consider the functor Then the map of spectra
is null-homotopic.
Note that is a special case of by Proposition 4.7. The following lemma is crucial for the proof of Theorem 5.1.
Lemma 5.3.
Let and Then the statement implies the statement
To prove this implication we use a general observation on the monads which are (non-deformed) tensor algebras.
Proposition 5.4.
Let be a dualizable category and let be a continuous exact endofunctor. Consider the monad put and denote by the left adjoint to the forgetful functor. Then for any the map is null-homotopic.
Proof.
It is convenient to prove a stronger statement on localizing motives. Namely, for we put It suffices to prove that the map
is null-homotopic in To see this, we first recall that by [E25c, Proofs of Theorem 8.1 and Proposition 8.2] the map
is null-homotopic. Thus, it suffices to show that the object is a retract of the object This follows from the standard argument: informally speaking, the non-negative grading on the tensor algebra corresponds to the multiplicative -action. More precisely, consider the map of monads on
which by adjunction corresponds to the composition
The map (5.2) has a left inverse induced by the map of -rings Passing to the categories of modules over the monads, we get the retraction
Passing to motives, we get a retraction
Now, the object is naturally a direct summand of Passing to reduced motives, we obtain the desired retraction
This proves the proposition. ∎
Proof of Lemma 5.3.
Suppose that holds. Let and be as in Put Consider the short exact sequence from Proposition 4.6, with modified notation:
We denote by the standard inclusions as in loc. cit., and denote by the right adjoint to Arguing as in the proof of Theorem 4.1, we obtain the isomorphisms
compatible with the transition maps. As in Proposition 4.6, we consider the functor right inverse to Recall that generates by construction, and by part (iii) of loc.cit., the adjunction counit is a split monomorphism, giving a direct sum decomposition Consider the morphism of monads corresponding by adjunction to the inclusion of as a direct summand. Put Then the maps of monads induce the chain of (strongly continuous exact) functors
By Proposition 4.7 and Proposition 2.22 (iii), and are naturally dualizable -categories and the functors and are -exact.
We claim that the functor is -coconnective. By Proposition 4.4 we need to check that the functor is left -exact. By Proposition 2.27 it suffices to show that the functor is left -exact. This is clear by construction: the latter functor is isomorphic to
and is left -exact by assumption.
For consider the cofiber sequence
Applying to we see that the map
is null-homotopic. Hence, the map
factors through Now the functor is left inverse to Applying Proposition 5.4 we see that the map
is null-homotopic. We conclude that the map
is null-homotopic. This proves the implication ∎
Before proving Theorem 5.1 we spell out a trivial sufficient condition for vanishing of a composition of maps between spectra equipped with finite filtrations. As usual, we denote by the totally ordered set for
Proposition 5.5.
Let and be integers, and let be a functor. Suppose that the following conditions hold.
for
The map
is null-homotopic for
Then the map is null-homotopic.
Proof.
We apply induction on The base is tautological.
We now prove a vanishing result which is in fact stronger than Theorem 5.1.
Proposition 5.6.
For the following statements hold.
the statement holds if
the statement holds if
Proof.
We apply induction on We already know that holds: this is exactly the assertion of Corollary 5.2.
Next, for we have the implication Indeed, suppose that holds and let be as in Then the quadruple satisfies the assumptions of so holds. The latter statement implies by Lemma 5.3. This proves that implies
It remains to prove assuming and for Let be as in We need to prove that holds. We may and will assume that and
Let be a -coconnective functor between dualizable -categories. We apply the construction from Proposition 4.8. By Propositions 4.9 and 4.7, it gives a direct sequence of dualizable -categories such that and each functor is -coconnective. Moreover, if are the transition functors for then generates and each monad is isomorphic to a deformed tensor algebra of some such that is left -exact.
Applying to the functor we see that Hence, we have an isomorphism
Thus, it suffices to prove that the map
is null-homotopic. We will apply Proposition 5.5 to the following functor:
We need to check that the assumptions of loc. cit. hold. Clearly, we have for It remains to see that for the map
is null-homotopic. This follows from Indeed, the functor is left -exact, the monad is the deformed tensor algebra of and we have
Applying Proposition 5.5, we obtain that the map (5.3) is null-homotopic, as required. This proves the proposition. ∎
Proof of Theorem 5.1.
The realization functor is -coconnective. It follows from Proposition 5.6 that for any and for any there exists such that holds. Namely, it suffices to take the smallest such that and put Therefore, for any we have
In other words, the map is an equivalence of spectra. This proves the theorem. ∎
6. Dévissage theorems for -theory and homotopy -theory
In this section we apply the results from Sections 4 and 5 to deduce dévissage theorems for coherently assembled abelian categories. One way to approach this is by generalizing the construction from [E25b, Theorem 0.3]. However, we choose a slightly different but similar argument.
Theorem 6.1.
Let be a coherently assembled abelian category. Let be a coherently assembled abelian full subcategory, such that the inclusion functor is strongly continuous, and generates via extensions and filtered colimits.
The map is an isomorphism for and a monomorphism for
The map is an equivalence of spectra.
Proof.
(i) The assumption implies that the induced functor is -coconnective. Hence, the statement of (i) is a special case of Theorem 4.1.
(ii) Consider more generally a -coconnective functor between dualizable -categories. We need to prove that the map is an equivalence. Arguing as in the proof of Theorem 4.1 we reduce to the special case when the monad is isomorphic to a deformed tensor algebra of some such that is left -exact. Arguing as in loc. cit. we find a dualizable -category and a -coconnective functor such that The latter spectrum is zero by Theorem 5.1 since we have by Proposition 2.22. ∎
We will see in the next section that in the situation of Theorem 6.1 the map does not have to be surjective, even when and are locally coherent and linear over a field.
Remark 6.2.
One can obtain a more precise version of part (ii) of Theorem 6.1 when every object of is an extension of two objects of Under this assumption a generalization of [E25b, Theorem 0.3] shows that the functor satisfies the following property: the monad is isomorphic to a deformed tensor algebra of some continuous exact functor such that is left -exact. By Proposition 5.6 for any and the map
is null-homotopic. Here () is the localizing invariant defined in (5.1).
7. Sharpness of the estimates
7.1. Estimates for theorems of the heart and dévissage
In this section we prove the sharpness of our estimates from Corollary 3.2 and Theorem 6.1 (i), even when we are dealing with small dg resp. abelian categories over a field.
Theorem 7.1.
Let be a field.
There exist small -linear abelian categories such that the inclusion functor is exact, each object of is an extension of two objects of and the map is not surjective.
For any there exists a -linear small -category such that the maps are isomorphisms for and and the map is not surjective.
We will need the following lemma, which is certainly known to experts, but we do not know a reference.
Lemma 7.2.
Let be an additive -category, such that for any object the ring is left or right artinian. Then we have for
Proof.
Since the functors commute with filtered colimits, it suffices to consider the case when is equivalent to the category of finite projective right modules over a connective -ring such that is (left or right) artinian. Let and Then by [BGT13, Theorem 9.53] and (for example) by the proof of [LT19, Corollary 3.5] we have where is the Jacobson radical. The ring is a finite product of matrix algebras over skew-fields, hence we may assume that is a skew-field. In this case all the spectra vanish for ∎
In the following proof we minimize the computations, focusing only on proving the sharpness results.
Proof of Theorem 7.1.
(i) Let be the cuspidal cubic curve over Denote by the category of vector bundles of finite rank on Denote by the dg category of acyclic bounded complexes of objects of By [Nee21, Lemma 1.2] this category has a bounded -structure whose heart is equivalent to the category of effaceable functors (see also Subsection 1.1, discussion before Proposition 1.3).
Denote by the abelian category of pairs where and is a nilpotent endomorphism. We have a fully faithful inclusion and we identify with its essential image. More generally, for denote by the (abelian) full subcategory of pairs such that Then the inclusion functors are exact and each object of is an extension of two objects of
We claim that for some the map is not an isomorphism, which would prove (i). It suffices to prove that the map is not an isomorphism.
By Proposition 2.25 (ii) we have so we need to show that this nil group is non-zero. By [Nee21, Proposition 2.4] the realization functor is an equivalence. Denoting by the category of with a split exact structure, we obtain a short exact sequence
By Lemma 7.2 we have Hence, The latter group is non-zero, since we have a surjection
This proves (i).
(ii) Take any pair as in (i) such that the map is not surjective. We will construct inductively the small -linear -categories with required properties, so that for all
First we construct We use the construction from [E25b, Theorem 0.3] and recall the assertions of loc. cit. We denote by by the quasi-abelian category of pairs where and is a subobject such that both and are in The forgetful functor induces a quotient functor on bounded derived categories. We define to be its kernel, so we have a short exact sequence
We have a semi-orthogonal decomposition Here the fully faithful functors are induced by the exact functors which are given respectively by The category has a bounded -structure, and we have an equivalence given by
Now, the cofiber sequence of -theory spectra for (7.1) is of the form
Both functors and are isomorphic to the derived functor of the inclusion Hence, by assumption the map is not surjective. In particular, the map is not a monomorphism, where is the right adjoint to But the functor is left inverse to the realization functor, hence the map is not surjective, as required.
Now suppose that for some we have constructed a -linear small -category with required properties such that the map is not surjective. We will construct the -category as follows. Put and consider the exact category with the induced exact structure. Consider another exact category from Proposition 3.7 with the same underlying category as (we use the small category version of loc. cit.). Recall from loc. cit. that a morphism in is an inclusion if both and are monomorphisms. We define via the short exact sequence
Our assumption on implies that the realization functor is an equivalence: this is Proposition 2.19 applied to Hence, by Proposition 3.7 we have a semi-orthogonal decomposition Here the functors and are induced by the exact functors given by resp. The category has a bounded -structure and we have an equivalence Moreover, by loc. cit. the maps are isomorphisms for
Now the cofiber sequence of -theory spectra for (7.2) is of the form
Both and are isomorphic to the realization functor Now, by Theorem 3.1 we have an isomorphism Hence, our assumption on implies that the map is not surjective. Hence, the map is not a monomorphism, where is right adjoint to The functor is left inverse to the realization functor, hence the map is not surjective, as required. This concludes the inductive step of the construction and proves the theorem. ∎
7.2. Estimates for higher nil groups of abelian categories
We give a closer look at the abelian category from the proof of Theorem 7.1 in the characteristic zero case. We recall the localizing invariant For a small abelian category we put We do the following computation, in particular proving the sharpness of the estimates from Corollary 4.12.
Theorem 7.3.
Let be a field of characteristic zero. Let be the cuspidal cubic curve over and let be the category of vector bundles of finite rank on Let be the (small) abelian category of effaceable functors
For we have
(7.3) In particular, we have
Moreover, we have
We elaborate on the meaning of the right hand side of (7.3). First, by we mean the negative cyclic homology over considered as a spectrum. The superscript “red” means the reduced part in the simplicial sense, where we consider the assignment as a functor see for example [E25c, Section 8].
For the proof of Theorem 7.3 we will need a certain short exact sequence, which might be known to some experts. We recall the notion of a lax pullback: if and are functors between small -categories, then is the category of triples where and If are stable and are exact, then the category is also stable, and it has a semi-orthogonal decomposition where the inclusions are given by resp. The same applies to -linear stable categories and -linear exact functors.
Proposition 7.4.
Let be as in Theorem 7.3. We identify the normalization with with coordinate We put and consider the pullback functors Consider the semi-free (associative) dg algebra where Then we have a short exact sequence in
Proof.
The fully faithful functor in (7.5) is the familiar categorical resolution of singularities from [KL15]. It is induced by the pullback functors: to the singular point and to the normalization
Put We need to compute the quotient First take the affine chart Then is a quotient of and induces a fully faithful functor
We have an equivalence
We consider the objects
Clearly, and together and generate (as an idempotent-complete stable subcategory). Hence, it suffices to compute the endomorphism dg algebra of the image of in the quotient of by (the stable subcategory generated by) .
First, the dg algebra is discrete, i.e. its homology vanishes in non-zero degrees. Hence, we may identify it with its Putting we have
Denote by the image of in By [Dr04] the dg algebra is quasi-isomorphic to
where has degree the multiplication is straightforward and the differential is given by Note that the cokernel of is identified with the shifted non-reduced bar complex computing the derived tensor product Since the non-reduced bar complex is quasi-isomorphic to the reduced bar complex, we have a quasi-isomorphism
This proves the proposition. ∎
We recall from [Ke64] that for an ordinary small additive category its Jacobson radical is a sub-bifunctor
We denote by the category with the same objects as in where the morphisms are given by
We will need a simple observation.
Proposition 7.5.
Let be a field and suppose that is an ordinary idempotent-complete -linear small category with finite-dimensional morphisms. Then we have for
Proof.
Indeed, the category is equivalent to a direct sum of (a possibly infinite collection of) categories of finite-dimensional vector spaces over skew-fields. For any skew-field we have for ∎
Now Theorem 7.3 is easily deduced from the following lemma. We first clarify the notation. For a -linear stable category (or for a -algebra, or for a scheme over ) we denote by the Hochschild homology over and similarly for and We will also use the notation for the higher nil versions, so for we have
We also note that for the -module is isomorphic to a direct sum of shifts of which implies an isomorphism
Lemma 7.6.
Let and be as in Theorem 7.3. As in Proposition 7.4 we identify the normalization with and denote by the category of vector bundles on of finite rank.
Consider the universal finitary localizing invariant of -linear idempotent-complete stable categories We have an isomorphism This gives a natural commutative square in the category
(7.6) -
Denote by the total cofiber of (7.6), considered as a complex of -vector spaces (forgetting the -action). Then for we have a natural isomorphism
Proof.
7.6 By Grothendieck’s theorem every vector bundle on is a finite direct sum of line bundles Hence, as a -linear dg category has a -indexed full exceptional collection Now, the Jacobson radical of is generated by the morphisms for Hence, the category has a (completely orthogonal) -indexed full exceptional collection formed by the images of It follows that the functor induced an isomorphism on In particular, the map is an isomorphism. The square (7.6) is obtained from the diagram
7.7 Let be an integer. Recall that by [Nee21, Proposition 2.4] we have an equivalence Hence, we have a cofiber sequence of spectra
By Proposition 7.5 and by the Dundas-Goodwillie-McCarthy theorem [DGM13, Theorem 7.0.0.2] we have the isomorphisms
Next, recall the semi-free dg -algebra from Proposition 7.4: it is given by where and Applying loc. cit. and [DGM13, Theorem 7.0.0.2] again we obtain
Combining (7.9) and (7.10) together with the cofiber sequence (7.8) we obtain the isomorphism (7.7). ∎
Proof of Theorem 7.1.
(i) Consider the complex of -vector spaces from Lemma 7.6. By loc. cit. it suffices to prove that Clearly, we have Next, by Proposition 7.4 for the dg - algebra from loc. cit. we have
Moreover, of the latter complex is isomorphic to Hence, we have and Applying Lemma 7.6 for we obtain
By Corollary 4.12 we have We conclude that and by the above we obtain This proves (i).
Remark 7.7.
It follows easily from the [E25c, Theorem 9.3] that for any the spectrum is naturally a -module, where has cohomological degree (recall that ). Moreover, it follows from [E25c, Theorem 9.1] that is acting locally nilpotently on since by loc. cit. we have The proof of Theorem 7.3 shows that for the abelian category we have an isomorphism of -modules.
8. Examples of compactly assembled -structures and coherently assembled abelian categories
8.1. Sheaves on locally compact Hausdorff spaces
For simplicity we consider only the continuously bounded -structures, and we only consider the categories of sheaves with values in a fixed category (not in a presheaf of categories like [E24, Sections 6.2-6.4]).
Proposition 8.1.
Let be a compact Hausdorff space and let be a dualizable -category.
There is a compactly assembled continuously bounded -structure on the dualizable category of presheaves such that
There is a unique compactly assembled continuously bounded -structure on the category such that the sheafification functor is -exact. Moreover, we have
If is a coherently assembled abelian category, then the category is also coherently assembled.
Proof.
(ii) It is well-known that there is a unique -structure on such that the sheafification functor is -exact. Moreover, this -structure is compatible with filtered colimits. To finish the proof, by Proposition 2.12 it suffices to construct a functor right inverse to the sheafification functor, such that is continuous, exact and left -exact. For a closed subset denote by the inclusion. We put
Then commutes with filtered colimits since so do and For the same reason is left -exact. It remains to show that is right inverse to the sheafification.
To see this, consider another (accessible exact) functor given by
We have natural transformations and where is the inclusion functor. By construction, for and for any compact we have
Hence, the maps and are isomorphisms of sheaves, as required.
Corollary 8.2.
There exist coherently assembled abelian categories for which the group is non-zero. For example, let be a field, and take Then is coherently assembled and
8.2. Nuclear modules over
Consider the abelian category of light solid modules over from [CS20, CS24]. It is generated by a single compact projective object We denote by the original version of nuclear solid modules over whose description will be recalled in the proof of the next proposition. This category is known to be dualizable (but not compactly generated). Recall from [E25a] that we have a bigger dualizable category such that the inclusion functor is strongly continuous and we have
Here the inverse limit is taken in Recall that we have
The latter category has a standard -structure since is regular. Our goal in the next proposition is to show that this -structure extends to a compactly assembled continuously bounded -structure on One can show by a more difficult argument that it further extends to a compactly assembled -structure on with the same heart, but the latter -structure is not continuously bounded.
The first part of the following proposition is known but we include the proof for completeness.
Proposition 8.3.
The category is locally coherent. The abelian category has homological dimension and we have
The standard -structure on induces a compactly assembled continuously bounded -structure on Moreover, the realization functor is an equivalence.
Proof.
(i) Consider the compact projective generator as above. Putting we have an equivalence Recall that the morphisms are simply the continuous maps of topological abelian groups It follows that we have an isomorphism Recall that we have an equivalence sending to Hence, we have an isomorphism Now, the object is injective in the small abelian category Moreover, any object of can be embedded into It follows that the ring is right coherent and we have equivalences In particular, these categories have homological dimension Since also has a compact projective generator, the equivalence follows for example from [Lur18, Proposition C.5.8.12].
(ii) Within the above notation, denote by the ideal of trace-class maps in the sense of [CS20, Definition 13.11]. Under the above identification the ideal consists of compact operators, i.e. maps such that for each the map has finite rank. We have a short exact sequence
By (i), the standard -structure on is compactly assembled and continuously bounded. By Corollary 2.13 it suffices to prove that the functor is -exact, where is right adjoint to
We have Hence, we need to prove that is flat as a left -module. This is known, but we give a proof for completeness. It is convenient to deal with the opposite algebra, i.e. with endomorphisms of We consider the poset
with reverse order: if for Then is directed. For each such consider the operator
Here is the topological basis of It is easy to see that as a right ideal in is given by the directed union:
where runs through functions as above. Now, each (as an operator) has zero kernel, and for we have It follows that is flat as a right -module, i.e. is flat as a left -module. This proves the proposition. ∎
8.3. Chase criterion
In this subsection we consider abelian categories in the setting of the following proposition. In fact they are exactly the dualizable -modules in as shown in [LLS]. They are also exactly the categories described by Roos’ theorem [Roo65]. Namely, these are categories of the form where is an associative unital ring and is an ideal such that
Proposition 8.4.
Let be a compactly assembled additive ordinary category. Consider the category of continuous additive functors. Then is a Grothendieck abelian category, satisfying (AB6) and (AB4*). Moreover, we have an equivalence
Here the target is the category of continuous exact functors.
Proof.
If is compactly generated, then and the stated axioms hold. The equivalence (8.1) is essentially a reformulation of Lazard’s theorem. Indeed, we have an equivalence and the full subcategory of flat functors is identified with
Now consider the general case. Restriction to gives a fully faithful functor Its right adjoint is induced by the functor Explicitly, for with we have
In particular, is exact and commutes with filtered colimits. Since (AB5), (AB6) and (AB4*) hold in they also hold in It is also clear that has a generator, namely where is a generator in
Finally, note that by the above observations the functor (8.1) is a retract of the functor
which is an equivalence by Lazard’s theorem. ∎
The following result is a generalization of Chase criterion of left coherence of an associative unital ring For completeness we give essentially a self-contained proof, which is purely categorical and covers the original theorem [Ch60, Theorem 2.1]. More precisely, the result in loc. cit. corresponds to the case when is the (compactly generated) category of flat right -modules and is the abelian category of left -modules.
Theorem 8.5.
Let be a compactly assembled additive ordinary category, and let The following are equivalent.
is coherently assembled.
has infinite products.
Proof.
. By Proposition 8.4 we need to show that the category has infinite products, provided that is coherently assembled. Restriction to defines a fully faithful functor
Now, exactness of implies that has a right adjoint, given by
Since the category of abelian groups satisfies (AB4*) (exactness of infinite products), the category has infinite products. Hence, so does the category as required.
We first consider the case when is compactly generated, and we identify Choose a sufficiently large regular cardinal with the following properties: the set of isomorphism classes of objects in is -small, and for any the group is -small. Then the category is abelian. Moreover, this small category has a collection of projective generators. More precisely, let be the full (additive) subcategory of objects of the form where is a -small set, and for Then we have an equivalence We identify with a full subcategory of via
Now, the colimit functor is identified with the precomposition functor which is exact and commutes with all colimits. Thus, it suffices to prove that the left adjoint functor is exact. Indeed, this would imply that is coherently assembled since so is
Now, exactness of means that for any the functor is flat. But this holds by assumption: if then the product exists in which exactly means that the functor is flat. This proves that is coherently assembled, or equivalently locally coherent.
Now consider the general case. We first observe that the category has infinite products. Indeed, it suffices to show that for a collection of objects their product exists in This product is computed as follows: we first take the product of in (which exists by assumption), and then take its image under the functor
8.4. Concluding remarks
We covered only some very basic non-trivial examples of continuously bounded -structures and coherently assembled abelian categories. We briefly mention the following directions.
One can consider abelian categories of almost modules where is a discrete associative unital ring, and is a two-sided ideal such that It is interesting to figure out some basic sufficient conditions (for example, in the commutative case) which imply that is coherently assembled. If this is the case, the functor does not have to be an equivalence, and one can consider the continuous -theory as -theory of the pair When this means that is right coherent and we obtain the usual -theory of
One can consider the categories of nuclear modules for more general adic spaces, in particular for formal schemes. For example, if is a regular noetherian commutative ring with an ideal we expect that the category has a compactly assembled continuously bounded -structure. If is not regular, there should still be a continuously bounded -structure with a coherently assembled heart, so one can consider -theory in this framework. We also expect nice -structures in the archimedean setting, say for nuclear gaseous modules over various gaseous rings.
One can construct a dualizable -category starting from a coherently assembled exact category Namely, denote by the same category with an exact structure from Proposition 1.32. Then we have a short exact sequence in
One can show that the category has a natural compactly assembled continuously bounded -structure, induced by the natural -structure on If is locally coherent, then this -structure is compactly generated, and the induced -structure on coincides with the (bounded) -structure from Proposition 1.3. In general the heart is equivalent to the category of sequential limit-preserving additive functors such that for any for any and for any compact morphism with there exists an exact projection in such that
Let be an associative unital ring, and consider the abelian category of right -modules. By Proposition 1.43, is coherently assembled if and only if is right coherent. In general, one can consider the left derived functors and define the non-coherence rank of to be
Suppose that One can show that the category is dualizable and we are in the situation of Question 0.7. One can define -theory of putting We are not aware of examples when is not compactly generated (under the assumption ).